Assumption Parish Operational Situation Summary Emergency Management: Parish

Assumption Parish Operational Situation Summary Emergency Management: Parish Emergency Declared 19 June 2012 Extended 30 days on 20 December 2013 Extended 30 days on 17 January 2014 Extended 30 days on 14 February 2014 Extended 30 days on 14 March 2014 Extended 30 days on 14 April 2014 Extended 30 days on 14 May 2014 Extended 30 days on 13 June 2014 Extended 30 days on 11 July 2014 Extended 30 days on 11 August 2014 Extended 30 days on 11 September 2014 Extended 30 days on 08 October 2014 Extended 30 days on 07 November 2014 State Emergency Declared 3 August 2012 LA Hwy 70 Open for traffic Mandatory Evacuation still in effect Security plans in place to secure sinkhole area Parish continues blog to inform residents Texas Brine issuing financial assistance for residents Bayou Corne Incident Command Center - HOTLINE 1-877-281-7311 Facilitating the [email protected] for questions APOEP and Sheriffs Office relocated Command Post to 1424 Hwy 70S As of 2/25/20 1 Assumption Parish Operational Situation Summary Public Safety:

As of 2/25/20 LSP providing air craft support LSP providing HAZMAT support to parish as needed LDWF providing air boat support to parish as needed No market or recreational hunting or fishing in restricted area LSA Command Post on site An office trailer has replaced the GOHSEP vehicle and will serve as a consolidated command post for the Assumption Unified Command All state agencies currently involved with the event will continue to be in attendance. GOHSEP has added a notification widget to their website indicating the current alert level for OG-1 and Hwy 70 2 Assumption Parish Operational Situation Summary Gas Volumes as of 12 Nov Well Operatio Reading Reading nal Well Date Time Status 11/11/20 ORW-01 14 9:03 AM shut in 11/11/20 ORW-02 14 9:09 AM shut in 11/11/20 Never ORW-03 14 Flared 11/11/20 ORW-04 14 8:17 AM shut in ORW-05 NA NA plugged 11/11/20 ORW-06

14 8:37 AM shut in ORW-07 NA NA plugged ORW-08 NA NA plugged 11/11/20 ORW-09 14 8:30 AM shut in 11/11/20 ORW-10 14 8:40 AM shut in 11/11/20 ORW-11 14 8:15 AM shut in ORW-12 NA NA plugged 11/11/20 ORW-13 14 12:03 PM shut in 11/11/20 ORW-14 14 7:56 AM open 11/11/20 ORW-15 14 9:07 AM shut in 11/11/20 ORW-16 14 8:34 AM shut in 11/11/20 ORW-17 14 8:46 AM shut in As of 2/25/20 Vent Well Flare Data (24-hr pd Ending 1 pm on 11/11/2014) Cumul Dishcha Gas Gas ative Previous Choke rge Static Diff. Diff. Rate Rate Gas Flared Flow Total Start End Time/ Size Orifice (PSIG (PSIG (Inch (MCF (MCF/ Flared (MCF SIWHP FWHP End Date Time Time Hours (1/64") (Inch)

) ) ) /H) D) (MCF) Total) (PSI) (PSI) Comments 1:00p 12:00 11/11/2014 m Pm 24 0 0.063 0.00 0.00 0.00 0.00 0.00 1,099 1,099 0.0 . 1:00p 12:00 11/11/2014 m Pm 24 0 0.063 0.00 0.00 0.00 0.00 0.00 1,680 1,680 0.0 . 11/11/2014 NA NA 1:00p 12:00 11/11/2014 m Pm NA NA NA 1:00p 12:00 11/11/2014 m Pm NA NA NA NA NA NA 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm NA NA NA 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00

11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm NA NA NA NA NA NA NA NA NA NA 24 NA 0 NA 0.063 0.00 0.00 0.00 0.00 0.00 1,076 NA NA NA NA NA NA NA 1,076 NA 29.5 NA NA . Plugged 1/27/2014. 24 NA

NA 0 NA NA 0.063 0.00 0.00 0.00 0.00 0.00 1,097 NA NA NA NA NA NA NA NA NA NA NA NA NA NA 1,097 NA NA 0.0 NA NA NA NA . Plugged 2/7/2014. Plugged 2/5/2014. 24 0 0.063 0.00 0.00 0.00 0.00 0.00 1,037 1,037 3.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 568 568 7.0

. 24 NA 0 NA 0.063 0.00 0.00 0.00 0.00 0.00 NA NA NA NA NA NA 112 NA 112 NA 0.0 NA 24 0 0.063 0.00 0.00 0.00 0.00 0.00 1 1 0.0 24 0.25 0.063 0.00 0.00 0.25 0.00 0.03 1,584 1,584 24 0 0.063 0.00 0.00 0.00 0.00 0.00 1,531 1,531 1.5 .

24 0 0.250 0.00 0.00 0.00 0.00 0.00 199 199 6.5 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 62 62 0.0 . NA . Plugged 7/15/2013. . 8.5 . 3 Assumption Parish Operational Situation Summary Gas Volumes as of 12 Nov 11/11/20 ORW-18 14 11/11/20 ORW-19 14 11/11/20 ORW-21 14 11/11/20 ORW-22 14

11/11/20 ORW-23 14 11/11/20 ORW-24 14 8:44 AM 8:42 AM 9:01 AM 8:59 AM 8:57 AM ORW-27 NA 11/11/20 ORW-26 14 8:22 AM 11/11/20 ORW-28 14 8:55 AM 11/11/20 ORW-29 14 8:52 AM 11/11/20 ORW-30 14 8:24 AM 11/11/20 ORW-31 14 8:20 AM 11/11/20 ORW-32 14 9:05 AM 11/11/20 ORW-33 14 11/11/20 ORW-36 14 12:29 PM 11/11/20 ORW-37 14 9:37 AM 11/11/20

ORW-38 14 9:14 AM 11/11/20 ORW-39 14 8:11 AM As of 2/25/20 1:00p 12:00 shut in 11/11/2014 m Pm 1:00p 12:00 shut in 11/11/2014 m Pm Never Flared 11/11/2014 NA NA 1:00p 12:00 shut in 11/11/2014 m Pm 1:00p 12:00 shut in 11/11/2014 m Pm 1:00p 12:00 shut in 11/11/2014 m Pm plugged shut in shut in shut in shut in shut in NA NA NA 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00

11/11/2014 m Pm shut in Never Flared 11/11/2014 NA NA 1:00p 12:00 open 11/11/2014 m Pm 1:00p 12:00 shut in 11/11/2014 m Pm 1:00p 12:00 shut in 11/11/2014 m Pm 1:00p 12:00 shut in 11/11/2014 m Pm 24 0 0.063 0.00 0.00 0.00 0.00 0.00 86 86 20.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 209 209 0.0 . NA NA NA

NA 0.0 24 0 0.063 0.00 0.00 0.00 0.00 0.00 1,063 1,063 2.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 110 110 2.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 298 298 16.5 NA NA NA NA NA .

Well tree valve closed 100% and locked 11-20-2013 24 0 0.063 0.00 0.00 0.00 0.00 0.00 173 173 5.0 . 24 0 0.625 0.00 0.00 0.00 0.00 0.00 28 28 15.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 59 59 27.0 . 24 0 0.625 0.00 0.00 0.00 0.00 0.00 42

42 2.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 13 13 21.5 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 90 90 14.0 . NA NA NA NA 24 6.5 24 0 0.063 0.00 0.00 0.00 0.00 0.00 247

247 1.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 1 1 0.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 105 105 15.0 . NA NA NA NA NA NA NA NA NA

NA NA NA NA NA NA NA NA 0.250 0.10 0.00 9.98 0.20 4.86 5,108 5,103 42.0 . 4 Assumption Parish Operational Situation Summary Gas Volumes as of 12 Nov 11/11/20 14 8:13 AM 11/11/20 ORW-41 14 8:10 AM 11/11/20 ORW-42 14 12:00 AM 11/11/20 ORW-43 14 8:02 AM 11/11/20 ORW-46 14 12:15 PM 11/11/20 ORW-48 14 12:07 PM 11/11/20 ORW-49 14

12:34 PM 11/11/20 ORW-50 14 12:37 PM 11/11/20 ORW-51 14 12:32 PM 11/11/20 ORW-52 14 12:17 PM 11/11/20 ORW-53 14 9:25 AM 11/11/20 ORW-54 14 12:10 PM 11/11/20 ORW-55 14 12:24 PM 11/11/20 ORW-56 14 12:20 PM 11/11/20 ORW-57 14 12:00 PM 11/11/20 ORW-58 14 9:40 AM OGRW- 11/11/20 01 14 7:59 AM ORW-40 As of 2/25/20 1:00p 12:00 shut in 11/11/2014 m Pm 1:00p 12:00 shut in 11/11/2014 m Pm NA shut in

open open open open open open shut in open open open open shut in shut in 11/11/2014 NA NA 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m Pm 1:00p 12:00 11/11/2014 m

Pm 1:00p 12:00 11/11/2014 m Pm 24 0 0.125 0.00 0.00 0.00 0.00 0.00 50 50 5.5 24 0 0.063 0.00 0.00 0.00 0.00 0.00 0 0 0.0 NA NA NA NA NA 24 0 0.063 0.00 0.00 0.00 0.00 0.00 13 13 24 1

0.063 0.00 0.00 24.63 0.02 0.47 760 760 24 6.5 24 1 24 0.75 25 NA NA NA NA NA 0.125 0.10 0.00 2.23 0.02 0.58 1,840 0.063 0.00 0.00 0.46 0.00 0.06 540 0.0 Install pump . Disconnected 12/5/2013. Now PMW 0.0 19S. . 33.0 . 1,840 5.0 long term pumping. Bucket Test 540 21.0 . 0.063 0.10 0.00 19.92 0.02 0.43 1,526

1,526 27.5 . 1.25 0.063 0.10 0.00 8.10 0.01 0.28 18.6 18.29 12.0 . 24 0.5 0.063 0.00 0.00 0.60 0.00 0.07 284 284 6.0 . 24 0 0.063 0.00 0.00 0.00 0.00 0.00 86 86 24 8.25 0.250 0.10 0.00 1.50 0.08 1.88 2,106 2,104 11.0 Long term pumping. Bucket Test 24 3 0.125 0.00 0.00 14.52 0.06 1.47 2,703 2,702

47.5 . 24 1 0.063 0.10 0.00 0.52 0.00 0.07 64 9.5 . 24 10 0.250 0.10 0.00 14.00 0.24 5.76 2,376 24 0 0.063 0.00 0.00 0.00 0.00 0.00 24 0 0.063 0.00 0.00 0.00 0.00 0.00 5,245 64 17 9.0 2,371 . 27.0 long term pumping. Bucket Test 17 0.0 . 5,245 2.5 .

5 Assumption Parish Operational Situation Summary Gas Volumes as of 12 Nov Well BC 2 OG 1 OG 2 OG 3A TBC 3 PMW 004m PMW 004s PMW 007m PMW 007s PMW 008m PMW 008s PMW 012m PMW 015m PMW 015s PMW 016m PMW 016s PMW 017m PMW 017s PMW 018m PMW 018s PMW 019s PMW 020m PMW 020s Time Data Collected 9:13 8:07 8:00 7:52 8:06 NA 9:31 NA 9:34 NM 8:32 NM NM 8:28 NM 7:50 NM 7:54 NM 9:23 8:50

NM 9:28 Tbg Pressure (psig) 4 327 216 706 2 NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA NA Csg Pressure (psig) 7 334 215 519 0 NM 46 NM 45.5 NM 0 NM NM 0 NM 55 NM 51.5 NM 3 10 NM 49

*Note: Cumulative gas flared for OGRW-1 represents combined gas flow from ORW-12 and TBC-RW-1 for all dates preceding March 4, 2013. Cumulative gas flared after March 4, 2013 represents only OGRW-1 (formerly TBC-RW-1). *Note: Total daily gas flared and cumulative gas flared is based on data through 12pm of previous day *Note: ORW-42 became a PMW on 10/2/13 ORW-42 is now PMW-019S *Note: PMW-012S became a ORW on 9/6/13 PMW-012S is now ORW-54 *Note: PMW-018 Flared for 50 MCF on 11-22-13 through flare 6 during the emergency response to mitigate gas kicks developed during well construction *Note: PMW-018 Flared for 1.53 MCF on 12-3-13 and 12-5-13 for 2.96 MCF for depressurization *Note: All well pressures are read daily including weekends and holidays with data submitted to conservationorder daily. 6 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. DEPT OF NATURAL RESOURCES - Conduct ventilation system inspections Daily berm inspections Survey rods on South berm (12/23/13) Survey settlement plates around containment berm (12/23/13) Water Level Survey (12/23/13) Survey around TBC tanks (12/23/13) Conduct daily well readings and flare maintenance Continue to depressurize ORW 54 Monitor bubble site 56 Logged and hard tagged at 4076 ft CB&I sampling Performed daily visual inspections of under slab vent systems and dewatered as necessary. Monitored cracks in south berm daily during berm inspections Began repairing and building up lowest area of south berm between R11 and R14 on Wednesday (1/1/14) Depressurize ORW 54 through separator so that wellhead can be replaced Downloaded transducer and track-it data from Saturday to Tuesday 12/28/14 to 1/1/14 Real-time air monitoring and alarm system for the ORWs and PMWs were operational as of Monday (12/30/13 Collected samples from outfall#2 on Monday (12/30/13) Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Site Code Status moved down to Code 2 on Saturday (01/04) Site Code Status moved down to Code 1 on Sunday (01/05) Performed general maintenance on Maurice Road: Pulled in side sloughing and graded road on Monday (01/06) Reinstalled survey rods in south berm that were lost during repairs on Monday (01/06). Rods will be surveyed twice per week on Mondays and Thursdays Installed new survey rods in north berm for surveying/monitoring on Monday (01/06). Rods will be surveyed once per week on Mondays. Continued to vent and flare ORW-54 through the gas/water separator with open choke Friday Monday (01/03 01/06) in an effort to reduce pressure low enough to change wellhead. Average well pressure: 10 12 psi. Performed daily routine flare maintenance Logged and tagged OG3A @3790 (1/3) 7 As of 2/25/20 Assumption Parish

Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Temporarily turned off unit at 1465 Sauce Piquante on Thursday (01/09) due to water flooding the blower intake unit from a broken water pipe on the house during the freezing weather. Unit will be turned back on when the water subsides. Conducted daily berm inspections. Began placing sand at the southwest corner of Pad 10 for construction of the new south berm on Wednesday (01/08), and will continue daily as weather permits. Collected weekly outfall samples on Tuesday (01/07) Measured water levels in MRAA wells on Tuesday (01/07) Conducted monthly sinkhole profile sampling on Wednesday (01/08) Conducted monthly bubble site monitoring on Wednesday (01/08) Conducted gas pressure monitoring of GP wells on Thursday (01/09) Collected water samples from OGRW-1 and ORW-54 on Thursday (01/09) Collected surface water sample from sinkhole (split with LDEQ) on Thursday (01/09) Performed geophone well maintenance on LA 10, 17, 19, and Geophone 1 on Thursday (01/09) Continued to vent and flare ORW-54 through the gas/water separator with open choke Tuesday Thursday (01/07 01/09) in an effort to reduce pressure low enough to change wellhead. Average well pressure: 11 12 psi. Dewatered ORWs 18 & 22 on Tuesday (01/07) Dewatered ORW 22 on Wednesday (01/08) Dewatered ORW 26 on Thursday (01/09) Performed daily routine flare maintenance Performed sinkhole profile sampling (1/8/14) Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Restarted vent system at 1465 Sauce Piquante on Friday (01/10) Removed air monitoring detectors from 170 Crawfish Stew on Monday (01/13) Conducted daily berm inspections. Continued placing sand for construction of the new south berm daily (except Monday due to weather), and will continue daily as weather permits. Current progress of the berm is from the southwest corner of Pad 10 to approximately 50 east of the Florida Gas pipeline (pending approval to cross pipeline) Conducted SWPPP inspection on Saturday (01/11) Collected monthly samples from industrial water wells on Friday (01/10) Instrumentation site general maintenance (battery checks, adjustments, etc) performed on Saturday and Sunday (01/11 01/12) As of 2/25/20 8 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Continued to vent and flare ORW-54 through the gas/water separator with open choke Friday Tuesday (01/10 01/13) in an effort to reduce pressure low enough to change wellhead. Average well pressure: 10 11 psi. A total of 3,872 gallons of water have been recovered from the separator. Shut in ORW-26 due to low pressure on Friday (01/10) Dewatered ORWs 19, 22, & 24 on Friday and Saturday (01/10 01/11) Shut in ORW-22 due to no pressure/gas flow on Sunday (01/12) Opened ORW-19 to flare on Sunday (01/12) Dewatered ORWs 6, 10, & 39 on Sunday (01/12) Dewatered ORWs 10, 19, 22, & 39 on Monday (01/13)

Performed daily routine flare maintenance Collected gas samples from pmw 16s, 18s, and OGRW-1 Logged OG-3A on Thursday (01/16). PRAL: 3,640. TD @ 3,646. Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of the new south berm daily, and will continue daily as weather permits. Current progress of placing the sand base for the berm is from the southwest corner of Pad 10 to approximately 135 west of the Florida Gas pipeline. Cleared and chipped trees on the Triche property for access road to ORWs 55 & 56 on Tuesday and Wednesday (01/14 01/15). Placed sand for access road on Triche property on Wednesday and Thursday (01/15 01/16). Collected monthly samples from MRAA wells Tuesday - Thursday (01/14 01/16) Surveyed rods in south and north berm and conducted weekly settlement plate survey on Tuesday (01/14) Collected weekly outfall samples on Tuesday (01/14) Completed sinkhole depth survey (fathometer) on Thursday (01/16). Depth: 185 190. Collected samples from OGRW-1, ORWs 6, 19, 21, 22, & 54, and DPVE wells 22 & 37 for outfall discharge parameters on Thursday (01/16) Relocated instrumentation site 33 away from the edge of the sinkhole on Thursday (01/16) Continued to vent and flare ORW-54 through the gas/water separator with open choke Tuesday Thursday (01/14 01/16) in an effort to reduce pressure low enough to change wellhead. Average well pressure: 10 11 psi. A total of 4,227 gallons of water have been recovered from the separator. Dewatered ORWs 6, 10, 19, 22, & 39 on Tuesday (01/4) Dewatered ORWs 10, 22, 32, & 39 on Wednesday (01/15) Dewatered ORWs 10, 19, 22, & 32 on Thursday (01/16) Drilled passive vent well PVW-03 to approximately 30 and set well (15 screen) on Thursday (01/16). Well will be grouted and completed today (01/17) Performed daily routine flare maintenance 9 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Sampled bubble site 23 and GPBS-23 (1/14/14) Conducted daily berm inspections. Continued placing sand for construction of the new south berm daily, and will continue daily as weather permits. Approximately 1,691 linear feet of the sand base for the berm has been completed. Cleared, chipped trees, and placed sand on the Triche property for access road to ORWs 55 & 56 Friday Monday (01/17 01/20). Cleared trees and placed sand for access road to ORW-57 Friday Monday (01/17 01/20). Approximately 317 of road completed. Completed monthly sampling event on MRAA wells Friday (01/17) Surveyed rods in south berm on Friday (01/17) Installed and surveyed settlement rods around ORW-38 on Friday and Monday (01/17 & 01/20) Completed sinkhole wagon wheel survey on Friday (01/17) Surveyed south berm, north berm, and settlement plates on Monday (01/20) Continued to vent and flare ORW-54 through the gas/water separator with open choke Friday Monday (01/17 01/20) in an effort to reduce pressure low enough to change wellhead. Average well pressure: 10 11 psi. A total of 4,717 gallons of water have been recovered from the separator. Dewatered ORWs 22 & 32 on Friday (01/17) Dewatered ORWs 22, 24, 6, & 18 on Saturday (01/18) Dewatered ORW 18 on Sunday and Monday (01/19 01/20) Grouted passive vent well PVW-03 on Friday (01/17) Drilled passive vent well PVW-01 Saturday and Sunday (01/18 01/19) TD @ 40 Grouted PVW-01 on Monday (01/20) Developed PVW-03 on Monday (01/20) Performed daily routine flare maintenance

Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of the new south berm daily, and will continue daily as weather permits. Approximately 2,334 linear feet of the sand base for the berm has been completed. Placed sand for access road to ORW-57 Tuesday Thursday (01/21 01/23). 834 feet of road completed. Length of road is now complete, however, will need to be built up for vehicular access. As of 2/25/20 10 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Collected weekly outfall samples on Tuesday (01/21) Measured water levels on MRAA wells on Tuesday (01/21) Surveyed settlement rods around ORW-38 on Wednesday (01/22) Collected samples for outfall discharge parameters from ORWs 21, 19, 6, 22, 54, DPVE 22 & 37, and OGRW-1 Conducted pressure monitoring for GP wells on Thursday (01/23) Continued to vent and flare ORW-54 through the gas/water separator with open choke Tuesday Thursday (01/21 01/23). On Thursday afternoon, the wellhead was changed out to become a permanent ORW. The separator will be removed and the well will continue to flare. A total of 4,937 gallons of water were recovered from the separator. Developed PVW-03 and PVW-04 on Tuesday (01/21) Installed Track-It on PVW-03 on Wednesday (01/22) Developed PVW-05 on Thursday (01/23) Drilled ORW-56, set well and bentonite seal on Tuesday (01/21). TD @ 140 Completed grouting ORW-56 on Wednesday (01/22) Drilled ORW-55 on Wednesday and Thursday, set well and bentonite seal. TD @ 135.5. Performed daily routine flare maintenance Performed logging of Oxy 1 Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of the new south berm daily, and will continue daily as weather permits. Approximately 2,354 linear feet of the sand base for the berm has been completed from the east side (Pad 10), and 70 from the west side (near LA-18). Placed sand to build up low areas on access road to ORW-57 on Friday (01/24) Surveyed the south berm and settlement rods around ORW-38 on Friday and Monday (01/24 & 01/27) Collected weekly outfall sample on Monday (01/27) Placed microbial bugs in MRAA wells on Monday (01/27) ORWs 18 & 56 were brought online to flare Friday (01/24) Grouted ORW-55 to the surface on Friday (01/24) Advanced and completed CPT-125 to 124.58 on Friday (01/24) Grouted inclinometer IPI-STR02 to the surface on Friday (01/24) Completed Geoprobe borehole for passive vent well PVW-02 on Friday (01/24) As of 2/25/20 11

Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Completed installation of meter run and connected ORW-56 to flare on Saturday (01/25) Advanced and completed CPT-126 to 136.98 on Saturday (01/25) Performed P&A activities on ORWs on Rig Road Saturday Monday (01/25 01/27) Installed transducers on ORWs 54 & 56 on Sunday (01/26) Constructed fences around PVWs on Sunday (01/26) and Monday (01/27) Rigged up and began drilling on ORW-57 on Monday (01/27) Performed daily routine flare maintenance Connected meter run and began flowing ORW 55 on Monday (1/27); starting pressure 47.5 psi; Choke set at 1; 1/8 orifice plate Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Earthwork was temporarily put on hold due to freezing weather conditions Tuesday and Wednesday (01/28 & 01/29). Continued placing sand for construction of the new south berm on Thursday (01/30), and will continue daily as weather permits. Approximately 2,354 linear feet of the sand base for the berm has been completed from the east side (Pad 10), and 70 from the west side (near LA-18). Collected samples for outfall discharge parameters from ORWs 21, 19, 6, 22, 54, DPVE 22 & 37, and OGRW-1 Conducted daily monitoring of passive vent wells Installed transducer in ORW-46 on Tuesday (01/28) In advance of the winter weather storm, the following ORWs and associated flares were shut in as a safety precaution on Tuesday (01/28): ORWs 52, 53, 55, 56, 48, 18, 4, 14, & OGRW-1 On Wednesday (01/29), the ORWs that were shut in due to the winter weather were brought back online. Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of the new south berm as weather permitted. A total of 3,011 linear feet of the sand base for the berm has been completed. Graded the north berm and main TBC access road on Monday (02/03) Surveyed the south berm, north berm, and settlement rods around ORW-38 on Friday and Monday (01/31 & 02/03) Conducted daily monitoring of passive vent wells Conducted daily well readings and flare maintenance Opened ORWs 19 & 26 to flare on Friday (01/31) As of 2/25/20 12 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Shut in ORW 26 on Monday (02/03) due to no flow Reinstalled transducer in ORW-41 on Monday (02/03) Flushed the screen in ORW-6, reinstalled the transducer, and restored the well to service on Monday (02/03) Prepared pumping equipment for ORW-38 pump test by testing connections, pumping capabilities, etc. and moved frac tank to location on Monday (02/03) Collected samples from ORW 55/56 (1/31/14) Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections.

Continued placing sand for construction of the new south berm as weather permitted. 3,122 linear feet of the sand base for the berm has been completed. Conducted SWPPP inspection on Tuesday (02/04) Placed sand on low areas on access road to ORW-57 on Thursday (02/06) Cleared path from Pad 9 to Pad 10 with marsh excavator in preparation of building containment berm on Thursday (02/06). Collected weekly outfall samples on Tuesday (02/04) Performed in-situ monitoring of industrial water wells and measured water levels on Tuesday (02/04) Performed monthly bubble site monitoring Tuesday and Wednesday (02/04 02/05) Surveyed the settlement rods around ORW-38 on Wednesday (02/05) Monitored gas pressures in GP wells on Thursday (02/06) Conducted daily monitoring of passive vent wells Conducted daily well readings and flare maintenance ORW-8 plugged: removed transducer, tubing, & flushed sand out on Tuesday (02/04) and completed grouting to the surface on Wednesday (02/05) ORW-38 pump test: Conducted pump and piping installations, connections and pump testing on Tues/Wed (02/04 02/05) and started pump test on Thurs (02/06). Completed gas collection piping connections at ORW-58 prior to drilling on Wednesday (02/05) Began drilling ORW-58 on Wednesday (02/05) and set well and seal on Thursday (02/06). TD @ 149. Screened 120 124. Began plugging activities at ORW-7 on Thursday (02/06). Removed tubing and flushed out sand. Completed building platform for water containment drum at bubble site 56 in order to remove temporary barge. Completed gas collection piping connections at ORW-57 on Thursday (02/06) Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of new south berm as weather permitted. A total of 3,356 linear feet of the sand base for the berm has been completed. Placed limestone on Triche property access road on Saturday (2/8) Conducted SWPPP inspection on Monday (02/10) 13 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Surveyed the south berm, north berm, and settlement rods around ORW-38 on Friday and Monday (02/07 and 02/10) Installed the casing for inclinometer IPI-STR01 to 85 on Saturday (02/08) and installed the inclinometer on Sunday (02/09) Conducted daily monitoring of passive vent wells Conducted daily well readings and flare maintenance ORW-7 and GP-ORW-8 was plugged and grouted to the surface on Friday (02/07) GP-ORW-5 was plugged and grouted to the surface on Sunday (02/09) ORW-38 pump test: Continued pump test Friday through Monday (02/07 02/10). The pump test was stopped on Monday (02/10) due to lack of gas production. Completed the installation of ORW-58 on Friday (02/07) by grouting the well annulus to the surface. ORW-57: drilled to 155.5 on Saturday (02/08), set well and seal on Sunday (02/09), and completed grouting to surface on Monday (02/10) ORW-58: installed meter run and Barton meter and began to flare on Saturday (02/08). Well development began on Monday (02/10). Installed survey rods around ORW-21 in preparation for dewatering test Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of the new south berm as weather permitted. A total of 3,526 linear feet of the sand base for the berm has been completed. Placed silt fence along access road on Triche property and access to ORW-57 on Wednesday and Thursday (02/12 02/13) Collected weekly outfall samples on Tuesday (02/11) Collected monthly samples from industrial water wells on Tuesday (02/11) Conducted in-situ monitoring of industrial water wells and measured water levels on Tuesday (02/11) Completed connections on inclinometer IPI-STR01 on Tuesday (02/11). Both IPI-STR01 & 02 are now online and recording. Surveyed the south berm and settlement rods around ORW-38 on Wednesday (02/12)

Conducted sinkhole depth survey (fathometer: 210) on Wednesday and Thursday (02/12 02/13) Conducted daily monitoring of passive vent wells Conducted daily well readings and flare maintenance Completed developing ORW-58 on Tuesday (02/11) A pressurized leak test was performed on the BC-2 wellhead assembly on Tuesday (02/11). No leaks were noted and the test pressure remained constant. Following the test, the Nitrogen gas that was used for the test was bled off and the well was placed back into monitoring status. Gas pressure on BC-2 has been 2-3 psi for the last two days. An additional soap and water bubble leak test was performed on Thursday. No leaks were found. As of 2/25/20 14 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Re-deployed AreaRaes for continuous air monitoring at PVWs 04 & 05 on Wednesday (02/12) Opened ORWs 57 & 58 to flare on Wednesday (02/12) Began forming surface completions at ORWs 55, 56, 57, & 58 on Wednesday (02/12) and completed pouring concrete pads on Thursday (02/13) Demobilized the barge and tank at bubble site 56 and replaced with a water collection drum placed on a wooden platform at the location on Thursday (02/13) Extended the risers on monitor wells for PVWs 01, 04, & 05 above accumulated water around the wells on Thursday (02/13) Laid out dewatering piping from ORW-14 to OGRW-1 for future dewatering activities on Thursday (02/13) Respec and CB&I split samples at PVW-03 CB&I collected gas samples from 57 and 58 Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of the new south berm as weather permitted. A total of 3,958 linear feet of the sand base for the berm has been completed. Placed silt fence along the access road to ORW-57 on Friday (02/14) Surveyed the old south berm, north berm, and settlement rods around ORWs 38 & 21 on Friday and Monday (02/14 & 02/17) Collected monthly surface water samples on Friday (02/14) Conducted daily monitoring of passive vent wells Conducted daily well readings and flare maintenance Laid out dewatering piping along ORWs on the south side of the highway for future ORW dewatering activities on Friday and Monday (02/14 & 02/17) Completed the fence enclosure around ORW-58 on Friday (02/14) Rechecked BC-2 wellhead assembly for leaks on Friday (02/14). No leaks were found. Tested equipment at ORW-21 in advance of the official startup of the upcoming ORW-21 pump test to verify pumping rates, connectivity to the MRAA, piping connections, etc. Friday Monday (02/14 02/17). Water level prior to pumping on Saturday was 5.77 below the top of the master valve measuring point, and achieved a drawdown of 46 at 1.1 gpm (pumped 280 gallons). Monday testing achieved a pumping rate of 2.0 gpm. Well pressure began with 7.0 psi and ended with 8.0 psi. CB&I collected a sample from PVW-03 for DEQ testing Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued placing sand for construction of the new south berm as weather permitted. The sand base layer of the berm has now been completely connected (4,363). As of 2/25/20 15 Assumption Parish

Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Surveyed and set grade stakes on the new south berm for sand base elevation on Tuesday (02/18) Collected weekly outfall samples on Tuesday (02/18) Conducted in-situ monitoring of the industrial water wells on Tuesday (02/18) Surveyed the old south berm, settlement rods around ORWs 38 & 21, and site settlement plate survey on Wednesday (02/19) Collected sinkhole profile samples on Wednesday (02/19) Conducted gas pressure monitoring at GP wells on Thursday (02/20) Conducted daily monitoring of passive vent wells Conducted daily well readings and flare maintenance Collected gas sample with CB&I from PVW-03 on Tuesday (02/18) Conducted preliminary work at ORW-21 and ORW-39 for upcoming pump test Tuesday Thursday (02/18 02/20). Replaced old pump with new electric pump, piping connections, etc. Continued laying out 2 HDPE dewatering piping along ORWs on the south side of the highway for future dewatering activities Tuesday Thursday (02/18 02/20). The majority of the piping has been laid out. Work continues on piping connections. Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Performed SWPPP inspection on Friday (02/21) Repaired washout areas resulting from rain events over the weekend on the new south berm on Monday (02/24) Surveyed the old south berm, north berm, and settlement rods around ORWs 38 & 21 on Friday and Monday (02/21 & 02/24) Conducted daily monitoring of passive vent wells Conducted daily well readings and flare maintenance Tested new pump at ORW-39 to assess well conditions on Friday and Saturday (02/21 02/22) Continued laying out 2 HDPE dewatering piping along ORWs on the south side of the highway for future dewatering activities Friday Monday (02/21 02/24). The piping leading to Outfall 001 (north berm) was completed on Monday. Conducted preliminary pump testing at ORW-30 to assess well condition on Saturday (02/22) Began pump test at ORW-21 on Monday (02/24). Initial pump rate of 1.0 gpm and 10 psi reduction in pressure transducer. Pumped 300 gallons. Began geotechnical borings on the new south berm on Monday (02/24). Completed two borings (20 & 21). CB&I collected a sample from Bubble Site #101 As of 2/25/20 16 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Continued geotechnical borings Tuesday and Wednesday (02/25 02/26). Completed B22, B23, & B24. All geotechnical borings on the new berm are now complete. Installed turbidity curtain along portions of the new south berm near Bayou Corne, near CPT-102, on Tuesday (02/25) and near CPT-103 on Thursday (02/27) Installed settlement plates along the new south berm on Thursday (02/27) Collected weekly outfall sample on Tuesday (02/25) Conducted in-situ monitoring of industrial water wells on Tuesday (02/25) Conducted the monthly MRAA sampling event Tuesday Thursday (02/25 02/27) Surveyed the old south berm, settlement rods around ORWs 38 & 21, and weekly site settlement plate survey on Wednesday (02/26) Conducted daily well readings and flare maintenance

Completed dewatering piping connections from ORWs 19, 18, 6, & 16 going to Outfall 003 (near ORW-9 west berm) on Tuesday (02/25) and placed a new 1,500gallon tank at Outfall 003 location on Wednesday (02/26). Conducted preliminary pump testing at ORWs 16, 19, & 18 to assess well conditions Tuesday - Thursday (02/25 02/27) Continued pump test at ORW-21 Tuesday - Thursday (02/25 02/27). Through day 4 of 30 planned days. Cumulative water pumped approximately 4,400 gallons. New flow meters were installed on Thursday (02/27). Performed well work over on ORW-4 Tuesday Thursday (02/25 02/27) to replace the well head and remove sand and silt in the well. Began work over on ORW-14 on Thursday (02/27). Flushed seal out. Work continuing today. Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Placed sand on low areas of the new south berm on Monday (03/03) Surveyed the old south berm and settlement rods around ORW-21 on Friday, Monday, and Wednesday (02/28, 03/03, and 03/05) Completed the monthly MRAA sampling event on Friday (02/28) Collected weekly outfall sample on Wednesday (03/05) Conducted in-situ monitoring of industrial water wells on Wednesday (03/05) Collected monthly samples from industrial water wells on Wednesday (03/05) Conducted monthly bubble site monitoring on Wednesday and Thursday (03/05 03/06) Conducted gas pressure monitoring at GP wells on Thursday (03/06) Conducted daily well readings and flare maintenance Conducted preliminary pump testing at ORWs 4, 6, 26, & 48 to assess well conditions over the last week (02/28 03/06) As of 2/25/20 17 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Continued pump test at ORW-21 daily. Completed 11 of 30 planned days. Cumulative water pumped approximately 19,400 gallons. Increased the pump rate to achieve 20 psi reduction in down-hole pressure on Monday (03/03) and 30 psi reduction on Wednesday (03/05). Performed well work over / redevelopment on ORWs 4, 10, 14, 16, and 37 over the last week (02/28 03/06). Shut in all PVW wells (1 5) on Friday (02/28) Reinstalled transducer in ORW-39 on Monday (03/03) Downloaded monthly transducer data from ORWs Monday Thursday (03/03 03/07) Set up equipment and prepared for dewatering test on ORW-14 Thursday (03/06) Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Stockpiled clay on Monday (03/10) Surveyed the old south berm and settlement rods around ORW-21 on Friday, (03/7) Began the monthly MRAA sampling on Friday (03/7) Conducted shut in test at NSDBS-056 on Friday (3/7) Collected Track-It data on monitoring wells on Friday (3/7) Replaced antenna at site 33 and repair communication issue on Friday (3/7) Conducted flyover for aerial photography of sinkhole on Monday (3/10) Conducted daily well readings and flare maintenance Began the 5-day feasibility pump test at ORW 14 on Friday (3/7) and at ORWs 4 and 30 on Saturday (3/11) Continued pump test at ORW-21 daily. Completed 15 of 30 planned days. Cumulative water pumped approximately 31,900 gallons. Increased the pump rate to achieve 40 psi reduction in down-hole pressure on Friday (3/7); producing a small amount of gas which later dissipated overnight; Pump rate was increased to 50 psi reduction during the day on Saturday (3/8) and decreased back to 40 psi that evening. Replaced filter pack in ORW 16 from 6/14 grade to a smaller grain size 16/30 filter pack.

Completed download of monthly transducer data from ORWs on Friday (3/7) Began equipment set up and preparation for dewatering test at ORW-26 on Monday (3/7) Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Stockpiled clay on pad 10 Installed overflow structure on new south berm Thursday (03/13) Began excavation of anchor trenches along the base of slopes of new berm on Thursday (3/13), in preparation for geotextile and GCL placement As of 2/25/20 18 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Surveyed the old south berm and settlement rods around ORW-21 on Wednesday (03/12) Completed the monthly MRAA sampling on Wednesday (03/12) Completed wagon wheel portion of Sinkhole Depth Survey Wednesday (3/12) Surveyed areas of existing and proposed water overflow structures on Thursday (3/13) Conducted daily well readings and flare maintenance Completed the 5-day feasibility pump test and transitioned into continuous pumping at ORW 14 on Wednesday (3/12) and ORW 4 on Thursday (3/13). Continued pump test at ORW-21. Completed 18 of 30 planned days. Cumulative water pumped approximately 44,00 gallons; An 8-hour test was conducted on Thursday (3/13) at maximum bottom hole pressure reduction for this test configuration; DNR/CBI were on site to monitor test. Depth of pump and pump rate were approximately 148.21 feet and 2.9 gpm, respectively. Redeveloped ORW 26 on Tuesday (3/11) and conducted 6-hour preliminary pump test on Wednesday (3/12). Began equipment set up and preparation for dewatering test at ORW 15 on Wednesday (3/12) Observation and analyze of BS #104 (outside of evacuation zone) (3/14/14) Performed daily visual inspections of Under Slab Vent Systems and dewatered as necessary. Conducted daily berm inspections. Total of 1,675 of geotextile/GCL deployed and covered with first lift of clay Completed Sinkhole Depth Survey on Friday (3/14); Deepest areas identified by fathometer was 230 and by wireline was 248. Completed routine survey of berms and ORW 21 settlement rods on Friday and Monday (3/14 and 3/17) Conducted daily well readings and flare maintenance Completed redevelopment of ORW 15 on Monday (3/17) Completed installation of water line ORW 57 to Outfall 2 frac tank on Friday (3/14) Restarted 5-day pump test at ORW 30 on Friday (3/14) Began 5-day dewatering test at ORW 26 on Sunday (3/16) OGRW 1 was shut-in on Friday (3/14) due to no pressure. Upon investigation, the pump was removed and well was tagged at 182 bgs on Monday (3/17). An electric pump will be reinstalled to decrease down hole pressure and increase pump rate. Continued pump test at ORW-21. Completed 22 of 30 planned days. Cumulative water pumped approximately 60,538 gallons; Average pump rate is 2.75 gpm Plugged and abandoned PVW-4 on Monday (3/17); small amount of gas encountered during retrieval of the auger; All appropriate safety protocols were followed; No issues. As of 2/25/20 19 Assumption Parish

Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Conducted daily berm inspections. Total of 2,950 ft of geotextile/GCL deployed and 2650 ft covered with first lift of clay Disked and dried clay placed along new south berm connector on Wednesday (3/19) Graded and compacted clay on the berm connector on Thursday (3/20) Completed survey of old south berm and ORW 21 settlement rods on Wednesday (3/19) Conducted daily well readings and flare maintenance Completed redevelopment of ORW 24 on Tuesday (3/18); Installed pump on Wednesday (3/19); Performed 6-hour aquifer step test on Thursday (3/20) Began work over of ORW 23 on Tuesday (3/18) and completed redevelopment of Wednesday (3/19) Completed 5-day feasibility pump test at ORW 30 on Wednesday (3/18) and installed permanent well head Continued monitoring of continuous dewatering of ORWs 4 and 14 Began work over at ORW 32 on Wednesday (3/18) and completed redevelopment on Thursday (3/20) Depressurized ORW-48 on Wednesday and Thursday (3/19-20) and installed pump and transducer on Thursday (3/20) Conducted 6-hour pre-test at ORW 15 on Wednesday (3/19); Redeveloped well and installed pump and transducer on Thursday (3/20) Plugged and abandoned PVWs 5, 1, and 2 on Tuesday (3/18), Wednesday (3/19), and Thursday (3/20), respectively. Monitored wells following P&A. No gas detected Began installation of electrical panel and connections associated with ORWs 57 and 48 on Thursday (3/20) in preparation for proposed continuous pumping. Pulled pump in OGRW 1 on Wednesday (3/19) and reinstalled in an effort increase pump rate. Continued pump test at ORW-21. Completed 24 of 30 planned days. Cumulative water pumped approximately 72,668 gallons; Pump rate is 2.84 gpm; Total gas vented 901scf; Continue 40 psi BHP reduction Removed Track-It data loggers from GP-37-1, GP-37-2, NSDMW017, GP-22-1, and GP-22-3 on Tuesday (3/18) Downloaded data and removed Track-It data logger from NSDMW006 on Wednesday (3/19) CB&I sampled Bubble Site 56 Conducted daily berm inspections. Total of 3,500 of geotextile/GCL deployed and covered with first lift of clay Disked and compacted clay placed along new south berm on Friday (3/21) Graded and compacted clay on the berm connector on Thursday (3/20) Completed survey of old south berm, north berm, and ORW 21 settlement rods Performed weekly TBC subsidence survey Reseal antenna at Site IPI-02 on Friday (3/21) Performed Sinkhole Profile Sampling (3/21) As of 2/25/20 20 Assumption Parish Operational Situation Summary - The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. Conducted daily well readings and flare maintenance Bled down and sampled TBC-3 on Friday (3/21) and installed PVC riser on well on Saturday (3/22) Installed privacy fence at ORW 48; installed pump; completed connections to the water discharge line to the frac tank on the north berm; and completed the service connections for power to ORW-48 on Friday (3/21) Performed a work over at ORW-1: flushed silt from the well and injected BMR on Friday (3/21) Developed ORW-1 and removed 1500 gal of water (3/24) Completed 5-day feasibility pump test at ORW 26 on Saturday (3/22)

Began work over 5-day feasibility pump test at ORW 24 on Friday (3/21) Installed pump and transducer in ORW-32 and completed connections to the water discharge line to the tank at ORW-15 on Friday (3/21); and installed Barton meter on Saturday (3/21) Performed 6-hour step test at ORW 32 on Saturday (3/22) and began 5-day feasibility test on Sunday (3/23) Began continuous pumping of ORW 48 on Saturday (3/22) Plugged and abandoned PVW 3 on Friday (3/21). Monitored PVWs 3 and 2 following P&A. Downloaded data and removed Track-It data logger from GP-5 and restored original valve Configuration on Friday (3/21) Continued pump test at ORW-21. Completed 28 of 30 planned days. Cumulative water pumped approximately 90,047 gallons; Pump rate is 2.96 gpm; Total gas vented 902scf; Continue 40 psi BHP reduction Continued monitoring of continuous dewatering of ORWs 4, 14, and 48 Conducted daily berm inspections. All GCL/geotextile for the new south berm alignment was covered with the first lift of clay on Wednesday (3/26). Total of 4,500 of geotextile/GCL was deployed Disked and compacted clay placed along new south berm on Tuesday and Wednesday (3/25-26) Began density/compaction and moisture testing on the berm extension between pads 9 and 10. All tests passed. Completed survey of old south berm and ORW 21 settlement rods Conducted daily well readings and flare maintenance Installed Barton meter and perform 6-hour step test at ORW-1 on Tuesday (3/25) and began 5-day feasibility test on Wednesday (3/26) Redeveloped ORW-2 on Tuesday (3/25), install infrastructure set up on Wednesday (3/26) and completed 6-hour step test on Thursday (3/27) Completed redevelopment of ORW 39 on Thursday (3/27) Installed air compressor near OGRW-1; Completed electrical connections and tested operation on Wednesday (3/26). Completed electrical connections to ORW-57 on Wednesday (3/26) Completed 5-day feasibility test at ORW-24 on Wednesday (3/26) Painted ballards and well heads on Tuesday and Wednesday (3/25 and 26) Installed meter run and choke valve at PMWs 18 and 16 on Thursday (3/27) to be used to vent pressure to the flare 21 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Completed air line connections from the new air compressor to OGRW-1 and ORW-14 on Thursday (3/27) and built enclosure for compressor. Relocated 6500-gallon tank to ORW-32, to be used in 5-day testing of ORW-1, 2 & 32 on Thursday (3/27) Completed pump test at ORW-21 on Tuesday (3/25). Cumulative water pumped approximately 98,565 gallons; Total gas vented 902scf; Continued monitoring of continuous dewatering of ORWs 4, 14, 48 and OGRW-1 Conducted daily berm inspections. Completed survey of old south berm and ORW 21 settlement rods on Friday (3/28) Conducted daily well readings and flare maintenance Flushed ORW-15 and assessed further development on Friday (3/28) Installed water lines from ORW-1, 2 & 32 to the 6500-gallon tank on Friday (3/28) Installed pump at OGRW-1 on Friday (3/28) Replaced existing pumps at ORWs 4 and 14 on Friday (3/28); Installed privacy panel at ORW 54 on Sunday (3/30) Continued monitoring of continuous dewatering of ORWs 4, 14, 48 and OGRW-1 Started 5 day Feasibility test at ORW -2 Conducted daily berm inspections. Compacted and graded south berm in preparation for compaction test on Monday 3/31) Stockpiled clay and increased clay in low areas on Tuesday (4/1) Completed survey of old south berm, north berm, and ORW 21 settlement rods on Monday Performed weekly TBC subsidence survey

Conducted daily well readings and flare maintenance Flushed ORW-15 on Monday (3/31) Completed development of ORWs 52 and 53 on Tuesday (4/1) Completed 5-day feasibility test at ORWs 32 and 1 on Monday (3/31) Completed 5-day feasibility test at ORW 2 on Tuesday (4/1) Begin 5-day feasibility test at ORW 39 on Tuesday (4/1) Downloaded Transducer data from wells on Tuesday (4/1) Continued monitoring of continuous dewatering of ORWs 4, 14, 48 and OGRW-1 Relocated Pump from ORW-1 to ORW-39 on Tuesday (4/1) in preparation of Pump Test As of 2/25/20 22 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Conducted daily berm inspections. Placed and compacted clay on New South Berm, emptying the soil stockpile on Thursday (4/3) Conducted six in-situ compaction tests and collected soil samples for laboratory geotechnical analyses on Thursday (4/3) Completed survey of old south berm, and ORW 21 settlement rods Performed initial survey of settlement plates along the New South Berm on Wednesday (4/2) Conducted daily well readings and flare maintenance Completed monthly downloading of pressure transducers ORWs and PMWs on Wednesday (4/2) Relocated pump from ORW-32 to ORW-54 on Wednesday (4/2) in preparation of pump test on Wednesday (4/2) Relocated pumping equipment from ORW-2 and installed at ORW-15 on Wednesday (4/2) Assemble custom wellhead and flange at ORW 57 and installed pump and tubing on Wednesday (4/2) Relocated 6,500-gallon water tank adjacent to ORW-54 on Wednesday (4/2) Completed well development of ORW-56 and reinstalled pressure transducers in ORW-52 and -56 on Wednesday (4/2) Installed power to ORW-54 and wired pump control panels into water tight pods for use at ORW-54 and 57 on Wednesday (4/2) Completed wiring of the 3 electric pump to be installed at ORW 57 on Thursday (4/3) Completed well redevelopment of ORW-54 on Thursday (4/3) and installed pump and two transducers Repaired the well head to casing connection at ORW-53 (4/3); permanent well head was re-attached and passed all pressure leak tests; Transducer re-installed. Continued monitoring of continuous dewatering of ORWs 4, 14, 48 and OGRW-1 Bubble site located ~1 mile north of bubble site 56 Conducted daily berm inspections. Conducted a total of 20 in-situ compaction tests and collected soil samples for laboratory geotechnical analyses on Friday and Saturday (4/4 and 4/5); all tests passed Sloped and compacted sides of berm and sealed the crest of the berm to enable rain runoff on Friday (4/4) Completed survey of old south berm, and ORW 21 settlement rods Surveyed centerline of the new south berm on Friday (4/4) Conducted daily well readings and flare maintenance Replaced wellhead and pump were removed from at ORW-54 on Friday (4/4). Pressure tested well head of ORW 53 and confirmed integrity on Saturday (4/5) Initiated continuous dewatering at ORW-54 on Saturday (4/5); Continued monitoring of continuous dewatering of ORWs 4, 14, 48 and OGRW-1 As of 2/25/20 23 Assumption Parish

Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Conducted daily berm inspections. Conducted SWPPP inspection on Tuesday (4/7) Graded gravel roads and pads on Tuesday (4/7) Conducted clay thickness investigation of areas along the south berm Tuesday (4/7) Completed survey of south berm, north berm and ORW 21 settlement rods and TBC subsidence survey Performed outfall and industrial well sampling Measured water levels of MRAA water wells Downloaded transducer data in MRAA water wells Sampled MRAA-08 for Gas-Water IFT and viscosity Conducted daily well readings and flare maintenance Replaced and existing pump and installed pneumatic pump in ORW 15 on Tuesday (4/7) Continued monitoring of continuous dewatering of ORWs 4, 14, 48, 54 and OGRW-1 Conducted daily berm inspections. Completed limestone placement on connector berm between pads 9 and 10 on Thursday (4/10) Placed clay on various low areas of new south berm on Thursday (4/10) in preparation for geotextile fabric and limestone placement Surveyed ORW 21 settlement rods Conducted daily well readings and flare maintenance Increased pump rate and performed 8-hour maximum draw down test at OGRW-1 on Thursday (4/10); pumping rate @ 3.4 gpm; maximum draw down was measured at 142.09 below the flaring; No gas produced. Completed 5-day feasibility test at ORW 39 on Wednesday (4/9); No gas produced Restarted 5-day feasibility test at ORW 15 on Wednesday (4/9); After 3 hours of pumping, the test was suspended due to the amount of suspended silt which caused a filter cake to form around the tubing screen. As a result, the existing pneumatic pump was removed on Thursday (4/10). The tubing was pressurized with nitrogen and the existing pump was replaced a smaller pump. Continued monitoring of continuous dewatering of ORWs 4, 14, 48, and 54. Completed subsidence monitoring as required as part of de-watering test at ORW-21 Performed log of OG3A and found a depth of 3779 Conducted daily berm inspections. Aerated and compacted berm in preparation for geotextile fabric placement on Friday (4/11) Completed 14 compaction tests and collected moisture/density samples from new south berm on Friday and Saturday (4/11-12); all tests passed Stockpiled limestone on Friday (4/11) 24 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Final survey of old south berm between R-10 to R-18 on Friday (4/11) Performed regular survey of the rest of old south berm on Friday (4/11) Conducted daily well readings and flare maintenance Performed leak test at TBC-3 on Friday (4/11); detected two minor leaks and corrected Pressurized ORW 15 with nitrogen on Saturday (4/12), due to its slow recharge rate Replaced pneumatic pump in ORW 14 with electric pump on Saturday (4/12) and attempted to perform an 8-hour drawdown test on Sunday (4/13); test was

interrupted due to pump failure. Finalized clay placement to elevation on new south berm (4/11) Conducted daily berm inspections. Continued placement of geotextile and limestone on new south berm on Monday (4/14) Stockpiled limestone on Monday and Tuesday (4/14 and 4/15) Collected samples from MRAA 1, 6, and 8 on Monday and Tuesday (4/14 and 4/15) Performed sinkhole depth survey on Tuesday (4/15) Collected outfall samples on Tuesday (4/15) Conducted daily well readings and flare maintenance Installed new transducer in OGRW-1 on Monday (4/14) Replaced pump in ORW 15 on Tuesday (4/15) and adjusted the flow to achieve water drawdown to approximately 30 feet below the well flange; Continued monitoring of continuous dewatering of ORWs 4, 48, and 54. Conducted daily berm inspections. Completed placement of limestone on new south berm and connector berm (between Pads 10 and 9) on Thursday (4/17) Continued sinkhole cleanup on Thursday (4/17) Conducted daily well readings and flare maintenance Began continuous pumping of ORW 15 on Wednesday (4/16); pump rate is 0.33 gpm; opened to flare Replaced electric pump in ORW 14 with a pneumatic pump and resumed continuous pumping on Thursday (4/17); average pump rate is 0.69 gpm; open to flare Continued monitoring of continuous dewatering of ORWs 4, 48, and 54. Conducted daily berm inspections. Conducted daily well readings and flare maintenance Removed the 1500-gallon tank from ORW-24; relocated to the storage area; removed the secondary containment berm on Friday (4/19) Continued monitoring of continuous dewatering of ORWs 4, 14, 15, 48, and 54. As of 2/25/20 25 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Conducted daily berm inspections. Removed water flow meters at Outfall 1 & 2; clean the screens and re-install on Monday (4/21) Began placement of riprap along designated slope areas of new south berm on Tuesday (4/22) Performed sinkhole cleanup on Monday and Tuesday (4/21 and 4/22) Measured pressure and flow rate to PVW-56 on Tuesday (4/22); static pressure at 0.8 psi, max flow rate measured at 3.4 scfh. Installed fitting pressure gauge at DPVE-23 on Tuesday (4/22); measured pressure at 0.5 psi Performed sinkhole profile on Wednesday (4/23) Collected outfall samples Conducted daily well readings and flare maintenance Completed preparations for pump at ORW 57 on Monday (4/21) Replaced pump at ORW-4 on Tuesday (4/22) Haul and stockpiled limestone on Monday (4/21) Continued monitoring of continuous dewatering of ORWs 4, 14, 15, 48, and 54. Conducted daily berm inspections. Conducted general ground maintenance around flares and wells: clearing weeds and marking pipe and valve locations on Wednesday (4/23) Completed riprap placement on new south berm on Wednesday (4/23) Performed sinkhole cleanup on Wednesday (4/23) Began stockpiling materials per contingency plan on Thursday (4/24) Preformed Sinkhole Profile (4/23)

Visited NSDBS-26 to record wellhead pressure and LEL levels at nearby monitoring and DPVE wells on Wednesday (4/23) Visited several bubble sites on Wednesday (4/23) to record water depth, obtain video of the current site conditions, and determine mitigation and monitoring feasibility; sites visited include NSDBS- 53, 02, 87, 09, 07, 58, and 98. Preformed Flare Sampling (4/24) Preformed Surface Water Sampling (4/24) Miller set elevations on new S berm (4/23) Conducted daily well readings and flare maintenance Removed pump and transducer from ORW 26 on Wednesday (4/23) Install transducer in ORW 9 on Wednesday (4/23) Redeveloped ORW 14 on Wednesday and Thursday (4/23 and 4/24) Continued monitoring of continuous dewatering of ORWs 4, 14, 15, 48, and 54. 26 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing - operational response activities. Conducted daily berm inspections. Continued stockpiling materials per contingency plan on Friday (4/25) Conducted daily well readings and flare maintenance Reinstalled pump in ORW 14 on Friday (4/25); opened to flare and resumed pumping Cleared grass/weeds surrounding Flare 2, wells ORW-36, 37, 49, 50 & 58 on Friday (4/25) Continued monitoring of continuous dewatering of ORWs 4, 14, 15, 48, and 54. Conducted daily berm inspections. Begin construction of the CPT/PMW pad off the south side of ORW 57 access road on Tuesday (4/29) Graded access road to ORW 54 on Tuesday (4/29) Placed additional riprap on south berm on Tuesday (4/29) Conducted SWPP inspection on Tuesday (4/29) Replaced desiccant and inspected data loggers at water level sites on Tuesday (4/29) Relocated instrumentation and inspected tilt meters at sites 22 and 26 on Tuesday (4/29) Performed routine weekly sampling of outfalls on Tuesday (4/29) Conducted daily well readings and flare maintenance Removed fuel tank, spill containment, etc. near OGRW-1 on Sunday (4/30) in preparation for planned TBC work-over of OXY-2 brine well in early May Redeveloped ORW-19 on Monday (4/28), using nitrogen flush Conducted preparations on Tuesday (4/29) for workover at ORW 6 Continued monitoring of continuous dewatering of ORWs 4, 14, 15, 48, and 54. Shut in ORW 56 (no gas)on (4/29) Conducted daily berm inspections. Completed construction of the CPT/PMW pad off the south side of ORW 57 access road on Wednesday (4/30) Placed (2) concrete barriers on the east side of OGRW-1 on Thursday (5/1), for protection during the OXY-2 well work-over Complete placing limestone at Rip-Rap location on S berm (5/1) Replaced desiccant and verified water levels of instrumentation sites 36 and 40 on Wednesday (4/30) Restored communication at site 60 by transitioning antenna, solar panel, and cutting brush on Wednesday (4/30) Installed new antenna and transmission line at site 33 on Wednesday (4/30) Cut trees between site 59 and 60 on Thursday (5/1) Relocated tilt meter at site 88 to stand alone post on Thursday (5/1). Performed quarterly maintenance at site 59 on Thursday (5/1) Performed maintenance at site water level site 35 (4/30) Installed flow gauge at GP-BS-26-3A (5/1)

As of 2/25/20 27 Assumption Parish Operational Situation Summary - The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. Conducted daily well readings and flare maintenance Completed a 1 water transfer line from ORW-15 to the tank at ORW-54 on Wednesday (4/30); removed the tank at ORW-15 Redeveloped ORW-6 on Wednesday (4/30), using nitrogen flush Excavated shallow buried pipelines under the access road east of OGRW-1 on Thursday (5/1); excavated a deeper trench, reconfigured the HDPE piping and backfilled; and re-graded the road, in preparation for OXY-2 well work-over Performed transducer and Track-it data downloads on Thursday (5/1) Continued monitoring of continuous dewatering of ORWs 4, 14, 15, 48, and 54. Increased pump flow at ORW 15 & 54 (4/30) Started ORW 56, 53, 18 back up and started flaring (5/1) Conducted daily berm inspections. Completed relocation of contractor trailer on Sunday (5/4) in preparation for planned Oxy 2 work over Switched tilt-meter at instrumentation site 88 and performed maintenance on IPI site 59 on Friday (5/2) Perform maintenance at instrumentation sites 99, 39, 50, 48, 44, 40, 01, 89 and IPI 01 on Friday (5/2) Cut limbs and perform maintenance at instrumentation sites 32, 34, 36 and 37 on Saturday (5/3) Completed spring maintenance on the tilt meters and water gauges at instrumentation sites 31, 29 and 54. Troubleshoot at 59 and return to working order on Sunday (5/4) Conducted daily well readings and flare maintenance Cleared grass around various flare, geoprobe , and ORW locations on Friday through Sunday (5/2 to 5/4) Continued monitoring of continuous dewatering of ORWs 4, 14, 15, 48, and 54. Conducted daily berm inspections. Monthly Bubble site inspections (5/6) Conducted daily well readings and flare maintenance Cleared grass around various flare, geoprobe , and ORWs (5/5 and 5/6) Continued monitoring and maintenance of continuous dewatering of ORWs 4, 15, 48, and 54. (5/5) Troubleshoot problem with thermocouple on ORW flare 4 (5/5) Bucket tests were done at the discharge points of ORWs 15, 48 and 54 (5/5) AAI assembled new grout hoses and tubing sets in preparation to push CPT 129 (5/5) Started work on OXY 2 (5/6) Removed QED AP2 pump at ORW 4 and replace it with a rebuilt pump and returned it to normal operating status. (5/6) Pushed cone at CPT 129 to refusal at 112.5ft bgs then the bore was grouted to surface (5/6) As of 2/25/20 28 Assumption Parish Operational Situation Summary - The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities.

Conducted daily berm inspections. Relocated board mats at CPT 129 in preparation for re-push (5/7) Completed hydromulching of new south berm (5/8) Cleared trees south of ORW-55 and built pad in preparation for CPT 131 (5/7 and 5/8) Installed gates at both ends of the new south berm (5/9) Placed mats in preparation for CPT 130, located on new south berm at IPT-STR-01 (5/9) Repaired silted fence and/or placement new fencing at various locations on Trische road, ORW 57 access road, and Rig Road (5/8) Conducted pressure monitoring of TBC shallow monitoring wells (5/8) Performed subsidence survey (5/7) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 15, 48, and 54. Completed thermocouple replacement on ORW flare 4 (5/7) Completed electrical service to the air compressor associated with ORW 14 (5/7) Advanced and completed CPT-134 at ORW 14; refusal at 136 bgs; grouted to ground surface without incident (5/7) Advanced and completed CPT-129R2 (located on ORW 57 access road); refusal at 133.3; grouted to ground surface without incident (5/8) Removed pump at ORW 14 (5/7) and began redevelopment/workover (5/8) Continued work on OXY 2 Conducted daily berm inspections. Completed final grading of new south berm (5/9) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 14, 15, 48, and 54. Completed workover at ORW 14 (5/9); Opened to flare and resumed pumping (5/10) Advanced and completed ORW 131; refusal at 117.2 bgs; grouted to ground surface without incident (5/9) Attempted to advanced CPT-130 at IPI-STR-01(5/10); due to electrical issues, the CPT cone was sent to manufacture for repairs; plan to advance and complete CPT on Wednesday (5/14) Performed weekly pressure transducer datalogger downloads from all Phase 3B ORWs (5/11) Continued work on OXY 2 Conducted daily berm inspections. Cleared weeds and overgrowth at ORW and geoprobe locations (5/12) Placeed board mats at proposed locations for CPT-127 and 128 on Sportsmans Drive (5/12) Placed board mats at CPT 130 (at IPI-STR-01 on new south berm) (5/14) 29 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Extended the pad at CPT-129 in preparation for building PMW-17 Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 14, 15, 48, and 54. Completed workover at ORW 4 (5/13) Advanced and completed CPT 127; refusal at 117 bgs (5/12) Performed transducer calibrations at PMW-08M & 08S; PMW-16M & 16S; PMW-12M & PMW-19S (5/13) Continued work on OXY 2 Conducted daily berm inspections. Placed board mats at proposed location for PMW-17 (access road for ORW 57) (5/14) Place hay wattles on Triche property as secondary measure for sediment retention (5/16) Place board mats on Triche property in preparation for installation of PMW-20 at ORW 55 (5/16)

Sampled MRAA wells 1, 6, and 8 (5/13-16) Completed placement of conduit and cable from LA 10 to meter at new surface site (located at the entrance of pad 3 access road) (5/15) Performed surface water sampling (5/15) Washed out and aligned sensors in LA 11 and LA 14 (5/12 14) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 14, 15, 48, and 54. Advanced and completed CPT 128 at 148 Sportsmans Drive; refusal at 140.6 bgs; grouted to ground surface with no incident (5/14) Advanced and completed CPT 130 (IPI-STR-01 on new south berm); refusal at 133.4 bgs; grouted to ground surface with no incident (5/15) Plugged and abandoned MRAA 03RM and 03RS (5/14), MRAA 03RD on (5/15) Filled Water Well 3 with additional grout to top of well casing (5/15) Completed borehole installation at PMW-17; TD 148.0 ft bgs; installed the nested well pair piping; cored and logged material (coring intervals 124to 134, 134to 144, and 144 to 149 no recovery below 140 ) (5/15); grouted and installed wellhead (5/16) Completed installation of new air compressor at ORW 4; connected to flare and pumping (5/16) Advanced and completed CPT 136 at ORW 37; refusal at 129.8; grouted to surface with no incident (5/16) Completed installation of new well piping in Oxy 2 (5/16) Place board mats at 148 Sportsmans in preparation for installation of PMW-07 at (5/19) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, and 54. Advanced and completed CPT 136 at ORW 37; refusal at 129.8 bgs; grouted to ground surface with no incident (5/16) As of 2/25/20 30 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Advanced and completed CPT 133 at ORW 4; refusal at 134.48 bgs; grouted to ground surface with no incident (5/17) Completed borehole installation at PMW-20; TD 129 ft bgs; installed the nested well pair piping; cored and logged material; coring intervals 104 to 114, 114to 118 (no recovery), 118to 126, and 126to 129 (no recovery); grouted (5/18) Cleared weeds and overgrown vegetation at well and flare locations (5/16) Prepared all hose, cable, transducer and electrical for pump for ORW-57 (5/16) Plugged and abandoned LA 15 and LA 16 on Oxy property (5/20) Collected weekly outfall samples (5/20) Surveyed new south berm and north berm settlement plates and CPT locations (5/20) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, and 54. Installed pump, transducer and permanent well head at ORW 57; connected to flare (5/19) Advanced and completed CPT 137 at ORW 58; refusal at 123.2 bgs; grouted to ground surface with no incident (5/19) Advanced and completed CPT 135 at ORW 15; refusal at 132.0 bgs; grouted to ground surface with no incident (5/20) Completed borehole installation at PMW-07; TD 130.5 ft bgs; installed the nested well pair piping; cored and logged material; obtained coring sample from 104 to 15 bgs (no recovered sample from 115 to 130.5 bgs); grouted (5/21) Cleared weeds and overgrown vegetation at well and flare locations Conducted daily berm inspections. Completed a SWPPP inspection (5/21) Completed mowing grass on the property north of Highway 70 (5/22) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, and 54.

Installed pressure gauges at PMW-17 & 20 (5/21) and at PMW-7 (5/22) Restored power to ORW 14 and restarted pump (5/21) Advanced and completed CPT 132 at OGRW 1; refusal at 132.0 bgs; grouted to ground surface with no incident (5/21) Change out nitrogen cylinder at flare 4; adjusted nitrogen flow at flares 2, 4 & 5 (5/22) Completed borehole installation at PMW-04; TD 130.5 ft bgs; installed the nested well pair piping; screened at 108 to 114 and 128.5 to 130.5; Attempted to get core samples starting at 102 bgs; drilled into something hard at the top of the MRAA; only able to obtain a core sample from 123 to 127 bgs; (5/22) will grout today (5/23) Logged OG 3A it tagged 3967 (5/22) Completed grouting of PMW 4 As of 2/25/20 31 Assumption Parish Operational Situation Summary - The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. Conducted daily berm inspections. Placed additional limestone to thin areas on the new south berm (5/27) Removed turbidity curtain from the west side of the original west berm, from the tie in of the new berm to the old south berm (~1000) (5/27) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, 54, 57. Installed pressure gauge at PMW-04S (5/23) Advanced and completed CPT 140; refusal at 131.5 bgs; grouted to ground surface with no incident (5/23) Completed development at PMW-20 (5/23) Installed surface completion at PMW-07 (5/23) Completed borehole installation at PMW-04; TD 130.5 ft bgs; installed the nested well pair piping; Attempted to core samples starting at 102 bgs; drilled into a hard object at the top of the MRAA; only able to obtain a core sample from 123 to 127 bgs; grouted (5/23) Perform transducer calibration at PMW-08M and ORW-38 (5/24) Advanced and completed CPT 138 at ORW 36; refusal at 114.2 bgs; grouted to ground surface with no incident (5/27) Monitored pressures at newly installed PMWs (5/27) Removed old boom and absorbent pads from sinkhole (5/27) Conducted daily berm inspections. Conducted SWPPP inspection; (5/28) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, 54, and 57. Replaced the valve and flushed the discharge line at ORW-4 (5/28) Changed out well head fittings at ORW-15, increased the size of the discharge hose (1/4 to ) (5/28) Mobilized drilling rig and equipment to PMW 15 on south berm; advanced to interim depth of 118 bgs (5/29) Completed development of PMW-07 M and installed well head fittings on both PMW-07M and PMW-07S (5/29) Installed well head fittings on PMW-20M and PMW-20S (5/29) Pushed CPT 139 to 127.5 ft (5/29) Placed and graded the connector berm from Pad-9 to Pad-10 (6/3) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, 54, 57. Removed and replaced pump at ORW 14; set pump at 145 below the well head flange (5/30) Removed and replaced pump at ORW 4; flushed discharge lines (5/30) Removed and replaced pump at ORW 15 (5/30) 32

As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Completed installation of PMW 15 on south berm on (5/31); collected and logged soil core samples (124 to 134, 134 to 144) Installed a pressure gauge at PMW-15-S (6/1) Assembled pipe and fittings for proposed water line for ORW-54 (6/1) Replaced track-its on ORW 48 and 54 (6/1) Completed development of PMW-15 (6/2) and removed board mats; painted bollards (6/3) Formed and poured surface completions and installed bollards at PMW 15 (6/1) Removed the well head and transducer at ORW-56; confirmed the physical measurement of the transducer cable and re-installed (6/2) Flushed discharge lines at ORW 4 and returned pump to operating status (6/2 and 6/3) Repaired discharge line at ORW 15 inside well casing and returned to operating status (6/2) Formed and poured surface completions and installed bollards at PMW 04 (6/2) Completed transducer & track-it downloads at 16 locations (6/3) Completed development of PMW-04 (6/3) Installed wellheads on PMW 04 and PMW 15 (6/3) Removed the well head, pump and transducer at ORW 57; repaired minor leak at the master valve and reassemble the pump, transducer and well head; returned the well to flowing condition (6/3) CB&I and Pisani complete bubble site filming run (6/3/14) Sampled the industrial water wells (6/3/14) Completed construction of containment berms at Outfall #1 and #2 and set new holding tanks in place (6/5) Placed concrete barriers on the old south berm: one located west of the MRAA pad and the second on the west end of the old south berm (6/5) Completed building access steps to the top of the new tank at outfall 1; installed a standing platform at the 1500-gallon tank for outfall 2 (6/10) Completed building up ORW 57 access road with limestone and additional sand (6/10) Set centerline survey rods on the new south berm alignment and surveyed the locations (6/4) Completed survey of remaining PMW wells and CPT locations (6/4) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, 54, and 57. Completed remaining transducer & track-it downloads (6/4) Installed pressure transducers and well heads at PMW-07 and PMW-04 (6/4) Calibrated the pressure transducer at ORW-38 (6/4) Flushed discharge line at ORW 4 (6/4, 6/5, and 6/8) As of 2/25/20 33 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Installed re-worked well head adaptors on each of 5 new PMW M wells (04, 07, 15, 17 & 20) (6/5) Changed pressure gauge at ORW-23 (6/6) Programed and installed pressure transducers at PMW-015 and PMW-017 (6/6) Completed the piping change over for the new tank associated with outfall #1 (6/8)

OG3A total depth measured at 3,875; casing seat measured at 1,959psi (6/9) Installed Track-it pressure data loggers on the 5 new PMW S wells; also installed desiccant holders at a number of locations (6/10) Placed and graded limestone near the storage tank at ORW-54; picked up old damaged board mats for disposal (6/11) Trenched across the ORW-57 access road in preparation for placement of a gas line for PMW-017 (6/12) Removed silt fence on the south section of the north berm, from the TBC access road to the old rig road, including the old sinkhole access road and the current sinkhole boat launch area; completed both sides of the old rig road from ORW-26 to the Geophone 3 pad. (6/11) Completed building up ORW 57 access road with limestone and additional sand (6/10) Cleaned frac tanks at outfalls #1 and #2 (6/18/14) Performed pressure monitoring of TBC and CBI shallow monitoring wells (6/12) Collected field measurements for under slab ventilation waters (6/17) Performed routine sampling from discharge outfalls (6/17) Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, 54, and 57. Removed pressure transducer from ORW-13; reprogramed and installed in PMW-16 M (6/11) Replaced pump at ORW-15 (6/11) Removed pressure transducer from ORW-38:, reprogramed and installed in PMW-20 M; installed a pressure transducer in PMW-20 S (6/12) Assembled pipes and fittings for gas line connection to respective flares for PMWs 017 and 20 (6/12) Moved 2 HDPE pipe into place for connection to ORW-49 (6/12) Completed gas line connection for PMW-020 to flare 6 at ORW-55 (6/13) Installed a galvanized tee, ball valve and union on the end of the meter run at PMW 49 in preparation for connection of a temporary gas line from PMW-007 (6/13) Installed a pressure transducer in PMW-17 M (6/16) Removed the pressure transducer and all well head fittings above the master valve at PMW-18 S (6/12); cleaned, applied new joint sealant, and re-constructed the well head fittings; measured and adjusted the location of the transducer; placed the transducer back into the well; test for leaks none found. (6/16) Performed 8-hour shut-in pressure test at ORW49; beginning well head pressure was 30.0 psi; pressure increased and stabilized at 47.0 psi; opened to flare and pressure decreased to 46.0 psi. (6/17) Removed pressure transducers at ORWs 11, 27, 29 & 39 (6/17) Completed venting of PMWs 016 S and 019 S; results confirmed connectivity. As of 2/25/20 34 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Completed sampling MRAA wells Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, 54, and 57. Completed venting of PMWs 007 S and 004 S; results confirmed connectivity (6/18) Removed existing pressure transducers at PMW-07 M and 04 M; re-installed pressure transducers and set to the same depth as the S wells (6/18) Performed 8-hour shut-in pressure test at ORW50; initial well head pressure was 31.0 psi; pressure increased and stabilized at 46.0 psi (6/18) Completed venting of PMWs 020 S; results confirmed connectivity. (6/19) Removed the existing pressure transducer from PMW-018 M; re-installed it and set it at the same depth as the S well (6/19) Attempted to conduct venting of PMW-018S (6/19); the well began producing water and the test was stopped; test was restarted on 6/20; terminated test due to water infiltrating the line, in addition due to the lack of pressure build up, the well will be re developed Completed gas line connections for PMW-017 (6/20) Successfully completed venting of PMW 019 S (6/20) Successfully completed venting of PMW 017 S (6/23) Successfully completed venting of PMW TBC-3 (6/24) Performed outfall sample at 1 and 2 SWPP inspection (6/26-27/14)

Started removing old silt fence (7/1/14) Removed turbidity curtain from N and W berms (7/1/14) Started Bubble Site survey (7/1/14) Remark the MPs on wells as needed (6/30 7/1) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 15, 48, 54, and 57. Logged OG 3 A (Hard Tagged at 3776 ft) (6/27/14) Installed electronic Flow Meter on DPVE 26 (6/27/14) Installed transducer in PMW -8M (6/27/14) Install new pump in ORW 4 (6/25/14) Redeveloped TBC 3 (7/1/14) (6.5 psi when work started, 12 psi at 0730 on (7/2) Started Monthly Transducer and track-it data downloads. (7/1/14) Leak test TBC 3 (6/27 7/1) Setup PMW venting equipment at TBC 3 and vented from 1055 hrs to 1102 hrs (6/25/14) As of 2/25/20 35 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing - operational response activities. Adjusted transducer at PMW 12M (6/25/14) Perform maintenance on Flare 2 and 3 (6/26/14) Shut in ORW 4, 14 and 15 (6/28/14) Opened ORW 52, 19 and 18 to flare (6/28/14) Leak tested TBC 3, PMW004s,007s,008s,015s,016s,017s,018s,019s,020s installed plugs. Performed monthly flare shutdown testing on all flares (6/30/14) Installed screens on the PMW M wells (ports open to atmosphere) (6/30/14) Remove transducer and cable at PMW 8 M and replaced failed cable. (6/30/14) Adjust depth of transducer in 12 M (6/30/14) Place additional limestone at ORW 57 (6/30/14) Remove silt fence on W side of Maurice Rd from ORW 9 to ORW 16 (6/30/14) Removed silt fence along east side of Maurice Road from ORW-9 to highway; removed fencing on Maurice Road from flare 5 to 100 west of ORW-6 (7/2) Completed bubble site monitoring (7/2) Completed pressure monitoring of geoprobe wells Performed routine outfall and industrial water well sampling (7/8) Recorded water levels of MRAA wells (7/8) Completed wagon wheel portion of the sinkhole depth survey (7/7) Collected water samples from several vent system (7/8) Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Download data from the flow meter on DPVE 26; removed the meter and piping (7/2) Re-developed BC-2 (7/2) Removed the QED pump and transducer at ORW 14 and performed work-over (7/2); Installed electric pump and transducer (7/4) Removed the transducer leads at ORWs 48 and 54 from the control module and reprogrammed the controllers (7/2) Redeveloped ORW-4; confirmed communication with the MRAA (7/3) Reworked the pump discharge lines of ORWs 48 and 54 (7/3) Conducted a 6-hr Step Rate Aquifer Test at ORW-14; ending at 1.1 gpm and drawdown of 39 feet; prelim investigation confirmed communication with the MRAA. Conducted slug test at PMW-016 M; preliminary results show communication with aquifer; test results will be analyzed and published under a separate header (7/5) Conducted slug tests at PMW-017 M and PMW-015 M; test results will be analyzed and published under a separate header (7/6)

Reconfigured the well heads at PMW-012 M and PMW-008 M and performed slug test on PMW 012M and 008M(7/7) Completed horizontal drilling under Hwy 70 (7/7) Installed 4 pressure transducers at MRAA-004 S & M and MRAA-005 S & M (7/8) 36 Complete connections from the pump at ORW-54 to the tank at Outfall #1 on the north berm (7/8) As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Performed weekly outfall sampling (7/15) Installed gas cap and flow meter at NSDBS-53; measured flow at 0.7 scfh (7/14) Cut small trees at NSDBS-23 in preparation for barge access (7/10) Installed digital gas flow meter, solar panel, and cell model housing at PVW-56 (7/10) Installed gas caps at NSDBS-2 (measured flow at 7 scfh) and NSDBS-58 (measured flow at 4 scfh) (7/15) Replaced cell modem at instrumentation site 59 (7/15) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 48, 54, and 57. Retrofit wellheads at NSDBS-26. Performed site reconnaissance at gas cap sites.(7/9) Installed electric service to ORW-4 (7/9);completed installation of pump and transducer in ORW-4; returned the well to continuous dewatering; preliminary flow rate was 0.75 gpm (7/10); pressure built up to 42 psi and opened to flare on (7/14); pressure is currently measured at 28.5 psi with drawdown measured at 21.54 Conducted slug test at PMW-007 M, PMW-004 M, PMW-018 M and PMW-020 M; results of test showed communication with aquifer (7/9) Cut small trees at NSDBS-23 in preparation of PVW installation (7/10) Logged OG1; ran gyro and sonar tools (7/10) Logged OG3A; Tag bottom at 3,648 (7/11) Installed and sampled PVW-26-6 (7/11) and PVW-26-10 (7/14) Downloaded weekly transducer and track-it data from wells undergoing continuous pumping (7/14) Excavated a trench next to ORW-26 for the gas piping from Flare 4 (7/17) Installed gas caps at NSDBS-7 and NSDBS-80; measured flow at 2scfh and 4scfh, respectively (7/16) Completed sampling of MRAA wells (7/17) Collected surface water sampling (7/16) Completed sinkhole profile sampling (7/17) Performed weekly outfall sampling (7/22) Measured and recorded water levels at MRAA wells (7/22) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 48, 54, and 57. Completed well installation at PVW-BS-23. (7/16) Performed monthly venting of PMWs 20S, 016S, and 017S. (7/16) Reconfigured gas piping north of HWY 70; (including gas lines for ORW-29, ORW-54, ORW-32, ORW-1 and PMW-19 S)(7/16) As of 2/25/20 37 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. -

Completed the gas piping south of the highway; Redirected all gas from north of Highway 70 to Flare 3; leak tested all connections: no leaks detected (7/17) Removed the QED pump and installed the sampling pump at ORW-40 (7/17); field monitored for water quality (pH, temp, conductivity, dissolved oxygen, oxidation reduction potential and turbidity) (7/19) Performed PMW venting of TBC-3 and PMWs 004 and 007 (7/18); Both PMWs exhibited communication with the aquifer; No flow was registered on the flow meter of TBC-3 therefore venting was suspended. (7/18) Installed sampling pump in ORW-39 and field monitored for water quality (7/20) Performed monthly venting of PMW-19S; venting was suspended due to low volumes of gas (7/20) Collected gas samples at PVWs and KGCs (7/30) Outfall sampling (8/5) Conducted monthly bubble site monitoring (8/5) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 48, 54, and 57. Continued monitoring of PVWs 05, 09, 23, and 56 Completed monthly ORW and PMW pressure transducer and track-it downloads (8/1) Installed digital pressure gauge at ORW 54 (8/1) Installed new electrical service at pad 9 for provide power to flare 2 and ORW-14 (8/5); resumed continuous pumping at ORW-14 (8/6) Reconfigured water discharge piping at outfalls #1 and #2 (8/5) Redeveloped PVW-05 (7/31 and 8/4-5) Redeveloped PVW-23 (8/4) Located bubble sites 109 and 110 Performed TBC facility subsidence survey (8/9) Surveyed berm rods, settlement plates and MRAA 004 and 005 (8/12) Performed monthly sampling of industrial water wells (8/12) Collect water level measurements and download data loggers from MRAA water wells (8/12) Outfall sampling (8/12) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 48, 54, and 57. Continued monitoring of PVWs 05, 09, 23, 47A and 47B, and 56 Continued water transfer from PVW-BS-09 and PVW-BS-23 to Outfall #2 As of 2/25/20 38 Assumption Parish Operational Situation Summary - The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. Redeveloped DPVE-26-3A and DPVE-26-1 (8/6) Rebuilt wellhead at PVW-47 (8/6) Removed digital flow meter from PVW-56 (8/11) Installed a meter run and the digital gas flow meter at ORW-21; open well to the flare (8/7); produced 16 cubic feet of gas in 20 minutes; Downloaded track-it and digital meter data removed meter and secured wellhead Flushed discharge line at Outfall 1 (8/11) Investigated ORW-14 pump issues; Resumed continuous pumping (8/12) Outfall sampling (8/19) Measured flow at NSDBS-02, 05, 07, 09, 23, 44, 53 and 58 (8/13) Completed quarterly site maintenance at sites (01, 18, 25, 26, 29, 41, 31, 32, 34, 37, 54, 59, and 89) to replace desiccant, verify calibration, cut trees to clear radio paths, and general site maintenance (8/15)

Replaced seals at instrumentation sites 01 and 34 (8/15) Placed pallets at instrumentation site 30 for access. Painted site 102. (8/14) Collected flare gas samples (8/19) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 14, 48, 54, and 57. Continued monitoring of PVWs 05, 09, 23, 47A and 47B, and 56 Continued water transfer from PVW-BS-09 and PVW-BS-23 to Outfall #2 Conducted communications test at ORWs, 24, 22, 32, 1, and 2 Redeveloped PMW 18S (8/18) Redeveloped PMW 19S and injected BMR (8/19 Modified wellhead risers at PVW-09 and PVW-23 (8/14) Redeveloped PVW-26-6 and PVW-26-10 (8/15) Replaced flow meters at PVW-26-6 and PVW-47B (8/15) Outfall sampling (8/25) Completed monthly sampling of MRAA water wells (8/21) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 48, 54, and 57. Continued monitoring of PVWs 05, 09, 23, 47A and 47B, and 56 Completed redevelopment of PMW 12M; Installed wellhead fittings (8/21); reinstalled transducer (8/25) Performed a 5-hour aquifer step at ORW 23 (8/20) 39 As of 2/25/20 Assumption Parish Operational Situation Summary - The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. Reinstalled the pressure transducer in PMW-19S (8/21) Performed 2-hour communication test at ORW 26 (8/21) Installed new Track-it data loggers on ORW-38 and BC-2 (8/22) Performed 2-hour communication test at ORW-30 (8/22) Attempted to perform 2-hour communication test at ORW-31; due to silting up of tubing, activities were suspended (8-23) Tagged the well tubing depth of ORWs 9 and 11 and OGRW-1 (8/26) Logged OG3A; TD 3,640 (8/26) Modified riser at PVW-09 with addition of 3-in riser (8/26) Installed KGC at new mud volcano (8/26) Installed KGC at NSDBS-111 (north of PVW-05) (8/25) Outfall sampling (9/2) Completed monthly bubble site monitoring (9/2) Installed transducers/data loggers in MRAA-01S, 01D, 06S, 08S, and 08D Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 48, 54, and 57. Pressure tested PMW 18S; Repaired nipple/valve connection fitting (8/27) Replaced transducer at PMW-008M (8/28) Tagged well tubing depths of ORWs 3, 13, 28, 29, and 37 (9/1) Completed directional drilling and installation of infrastructure in preparation of ORW 51 installation (9/2) Redeveloped ORW-31 (9/2) Outfall sampling (9/9) Downloaded transducer data and measured water levels in MRAA water wells (9/5)

Completed monthly industrial well sampling (9/9) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 48, 54, and 57. Installed electrical service and a secure meter panel for the area RAE located at ORW-54 (9/4) OG3A tagged (9/5); hard tagged at 3462 Completed MRAA evaluation test (9/5) Completed installation of ORW 51 (121 Sportsmans Drive); TD at 136.5 bgs and screened from 107.5 bgs to 111.5 bgs; installed pressure gauge (9/5/14) Removed one of two transducers from ORW-54 (to be installed in ORW 51); flushed out approximately 9 inches of silt; reinstalled pump and sealed well head; restarted pump (9/8) 40 As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - - Completed work over and well construction modifications at PMW-19S (9/8); Removed existing tubing and replaced with new tubing, screened at 115.7 bgs to 120.7 bgs (perforations at 117.0 bgs to 121 bgs); redeveloped using both water and nitrogen; communication established; installed track-it data logger and pressure gauge; no leaks were detected. Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 4, 48, 54, and 57. Completed PMW-19S well construction modifications (9/10); adjusted the depth and installed the transducer in the screened zone of the well Removed, adjusted and re-installed the transducer in PMW-12M to match the transducer depth in PMW-19S (9/10) Completed development of ORW 51 (9/11); Poured the concrete surface completion (9/10); installed choke valve and meter run; completed connections to the gas piping; Installed barton meter on ORW 51 (9/12); opened to flare; and installed fence Flushed silt from ORWs 37, 29, 28, 13, 11, 19, and 17 Install pump and controls at ORW-31 in preparation for 2-hour communications test (9/15) Reset transducer depth at ORW-15 (9/16) to 121.62 btmp Installed pump and controls at ORW-28 (9/16) in preparation for 5-hour aquifer step test Outfall sampling (9/23) Completed transducer downloads and water level measurements of MRAA water wells (9/23) Completed TBC quarterly Fenstermaker survey (9/24) Replaced Taurus seismographs in LA 10 and LA 18 (9/19) Conducted maintenance on Geophone 1 to address moisture and corrosion issue (9/19) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Completed 5-hour aquifer step test/communication test at ORW 28 (9/17) Completed 2-hour communication test at ORW 37 (9/18) Completed 2-hour communication test at ORW 9 (9/19) Completed work over at OG-2 (9/18) Logged OG-3A; tagged at 3472 (9/19) Completed 2-hour communication test at ORW 6 (9/20) Completed 5-hour aquifer step test/communication test at ORW 29 (9/23) Installed meter run, Barton meter, an a control valve to GOW-9-1; connected to flare; initial pressure was measured @ 47 psi; after flaring for 20 minutes pressure 41 reduced to 29 psi

As of 2/25/20 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Outfall sampling (9/23) Installed a new valve assembly and rerouted the riser at PVW-09 (9/29) Installed a pressure gauge at PVW-23 (9/29) Installed KGC at PVW-05 (9/30) Secured riser and reinstalled tank line at PVW-09 (9/30) Performed maintenance to LA-10; back on-line (10/1) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Completed 5-hour aquifer step test/communication test at ORW 11 (9/24) Completed P&A of ORW 27; a steel plate monument with well number and date was placed over the area of the abandoned well (9/24) Completed P&A of GOW-9-4 (9/24) Completed 5-hour aquifer step test/communication test at ORW 43 (9/25) Completed P&A of GOW-9-2 and GOW-9-5 (9/25) Completed modified 1-hour communication re-test at ORW 32 (9/26) Completed modified 1-hour communication re-test at ORW 24 (9/29) Completed modified 1-hour communication test at ORW 13 (9/30) Installed water/gas separator at GOW-9-1 (9/30) Completed 2-hour communication test at OGRW 1 (9/30) Conducted daily berm inspections. Placed and graded limestone at the ORW-54/PMW-19S area & at the PMW-20 area on the Triche property (10/1) Outfall sampling (10/7) Surveyed the ORW-51 location (10/2) Competed monitoring and collecting monthly pressure and flow measurements of KGCs and PVWs Performed site maintenance of various instrumentation sites. Completed housing installation at BS-56 (10/4) Installed KGC at NSDBS-3 (10/6) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Completed 5-hour aquifer step test/communication test at ORW 41 (10/1) Installed a storage tank at Outfall 3 (west berm, near ORW-9) (10/2) Completed monthly transducer downloads (10/3) OG3A was logged and tagged at 3440 (10/3) Completed a modified 1-hour communication/aquifer test ORW 23 (10/6) As of 2/25/20 42 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Conducted daily berm inspections

Reshaped containment at Outfall #1, built platform and installed and connected 150 gallon tank(10/13) Performed SWPPP inspection following Mondays rain event (10/13) Cut vegetation around ORWs, inclinometers and other instrumentation (10/14) Outfall sampling (10/14) Re-surveyed select points around TBC gate and entrance road (10/13) Competed quarterly bubble site monitoring (10/8) Began quarterly MRAA monitoring well sampling (10/13-14) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Completed 2-hour communication test at ORW 03 (10/13) Stabilized vent pipe on PVW-05 (10/9) Continued daily flaring of GOW-9-1 Completed quarterly MRAA sampling (10/17) Performed sinkhole depth survey (10/17) Completed flare sampling (10/17) Outfall sampling (10/21) Competed monitoring and collecting pressure and flow measurements of KGCs and PVWs (10/21) Replaced desiccant in automated flow system, install water level transducer (10/21) Collected monthly samples from industrial wells (10/21) Downloaded transducer data and manually measured water levels in MRAA water wells (10/21) Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Completed quarterly venting of PMWs 04, 07, 016, 017, and 020 (10/16-17) OG3A was logged and tagged at 3486 (10/17) Completed sinkhole profile sampling (10/22) Conducted daily berm inspections. Placed board mats at ORW 57 and ORW 43 in preparation of MRAA 09D and 06D installation Performed monthly gas pressure monitoring of GP wells (10/24) Routine weekly outfall sampling As of 2/25/20 43 Assumption Parish Operational Situation Summary The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. - Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Performed 1-hour gas depletion tests on ORWs 10, 22 and 16 (10/27-28) Collected bottom hole pressure and depth to water measurements for transducer verification at select wells Performed surface water sampling Conducted daily berm inspections. Performed lawn maintenance in areas associated with wells Sampled MRAA 6D (11/3) Performed monthly bubble site monitoring (11/4) Completed wagon wheel survey of sinkhole (11/5) Routine weekly outfall sampling Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57.

Performed monthly transducer downloads (10/30-31) Completed installation of MRAA 06D; obtained core samples; drilled to total depth of 340; screened from 315 to 325 Performed 1-hour gas depletion tests on ORWs 17, 18, and 19 (11/3-4) Replaced transducer at PMW-007S (11/4) Performed work over ORW-6 (11/3) Surveyed MRAA 02, MW-1 and MW-2 (11/5) Sampled industrial wells (11/4/14) Sampled MRAA 9D (11/10) Routine weekly outfall sampling Surveyed MRAA 6D and 9D Conducted daily well readings and flare maintenance Continued monitoring and maintenance of continuous dewatering of ORWs 48, 54, and 57. Logged OG3A and ran string shot; TD at 3,524 Installed discharge piping from ORW 40 to the tank at outfall 1 (11/6); provided well with electrical service (11/7) and began continuous dewatering (11/11) Completed installation of MRAA 09D; obtained core samples; drilled to total depth of 340; screened from 310 to 320 Performed 1-hour gas depletion test on ORW 58 (11/10) Calibrated and performed maintenance on seismic array system As of 2/25/20 44 Assumption Parish Operational Situation Summary DHH / Office of Public Health Section for Environmental Epidemiology and Toxicology Air Monitoring Sample Data SEET has received and is analyzing community ambient air sampled 5/2/14, 5/6-7/14, 5/9/14, 5/13-14/14, 5/16/14, 5/20-21/14, 5/23/14, 5/27/14, 5/30/14, 6/3-4/14, 6/6/14, 6/13/14, and 6/17/14 (MultiRAE) and 5/22-27/14 (MAML). SEET will issue a letter to the parish in reference to these findings once the review of the data has been completed. SEET has received and is analyzing sample results for air at bubble sites collected 5/7/14, 5/14/14, 5/21/14, 6/4/14, and 6/17/14 (MultiRAE). SEET will issue a letter to the parish in reference to these findings once the review of the data has been completed. As of 2/25/20 Assumption Parish Operational Situation Summary DHH / Office of Public Health Section for Environmental Epidemiology and Toxicology Industrial Water Well Sampling Data No new water well data received during this operational period. As of 2/25/20 46 Assumption Parish Operational Situation Summary DHH / Office of Public Health Section for Environmental Epidemiology and Toxicology Slurry Water Sampling Data

No new slurry data received during this operational period. As of 2/25/20 47 Assumption Parish Operational Situation Summary DHH / Office of Public Health Section for Environmental Epidemiology and Toxicology Surface Water at Bubble Sites No new surface water data received during this operational period. As of 2/25/20 48 Assumption Parish Operational Situation Summary Department of Environmental Quality Air monitoring for VOCs continues with no dangerous levels detected off site (1000 away). On going Water sampling/water quality monitoring in Bayou Corne and Grand Bayou indicates no water pollution associated with the incident to date. To date outdoor air monitoring has been conducted in response to the receipt of 97 rights of passage. 57 indoor air monitoring inspections have been conducted in response to citizen requests. Private residence air monitoring is available for scheduling upon citizen completion of the Right of passage form. Please call (225) 219-3015 to notify DEQ if you desire monitoring.. LDEQ Command Post on site. Mobile Air Monitoring Lab (MAML) scheduled to be on site on a recurring basis collecting continuous data on methane/ non-methane VOCs and H2S. DEQ is conducting community air monitoring DEQ has submitted comments on Texas Brines proposed water control structure to DNR Approved emergency water discharge permit allowing discharge of water back into the sinkhole (9/6) Has begun water sampling weekly in the Bayou near sinkhole water discharge

Collected Hydrocarbon samples within sinkhole Mobile Air Monitoring Lab (MAML) departed site 5/27/14 Held public hearing on TBC LPDES permit application (9/9/14) As of 2/25/20 49 Assumption Parish Operational Situation Summary Department of Transportation and Development Regular monitoring of roadways and bridges in the area for ground movement Developed a traffic contingency plan if needed As of 12 Nov 2014 DOTD monitoring equipment has not detected any permanent bridge or highway movements nor any noteworthy trends. No significant changes in the surface of Hwy 70 profiler runs In an abundance of caution DOTD crews are working to establish a fifth CORS site, which is currently planned to be established in the Bayou Corne community. This fifth CORS device was part of the original package, and is not the result of any new developments Held stakeholders meeting for Hwy 70 reroute options 15 May Performed Hwy 70 profiler run (10/21) As of 2/25/20 50 As of 2/25/20 51 Assumption Parish Operational Situation Summary Department of Transportation and Development (contd) ACTIVITY CORS 1 CORS 2 CORS 3 CORS 4 ACTIVITY SAA 1 SAA 2 SAA 3 SAA 4 SAA 5

SAA GENERAL SENSR 1 SENSR 2 SENSR 3 SENSR GENERAL STATUS/REMARKS Installed on Monday, April 1, 2013. No significant movements off of baseline noted Installed on Tuesday, July 17, 2013. No significant movements off of baseline noted Installed on Monday, April 8, 2013. No significant movements off of baseline noted Installed on Tuesday, April 9, 2013. No significant movements off of baseline noted Reports from the active CORS sites are published daily and provided online at the following FTP site: ftp://mimir.lsu.edu|anonymous:[email protected]:2123 interface account has been created. Credentials for accessing CORS911 sites via web site was distributed on April 29, 2013. Web Installed April 23, 2013. No significant movements off of baseline noted. Installed April 25, 2013. No significant movements off of baseline noted Installed July 2, 2013. No significant movements off of baseline noted. Installed April 30, 2013. No significant movements off of baseline noted. Installed April 24, 2013. No significant movements off of baseline noted. Data from each station is being downloaded to DOTD hourly with manual analysis performed periodically. Options to provide reliable cellular communications to each site are being investigated. Automated alerts are being sent to B. Fernandez and J. Rauser for situational awareness and action as needed. Installed all equipment in April 2013. Training is completed. No significant movements off of baseline noted SAP Contract No. 4400001862, State Project No. H.010571.1, LA 70 Bypass, Stage 0 Feasibility Study in separate document. Stage 0 meeting was held for possible bypass road on 27 MAR. A stakeholder meeting regarding the potential Hwy 70 bypass road was held on April 11, 2013. HWY 70 CONTINGENCY PLAN Developed August 11, 2012 by District 61 . EMERGENCY Draft plan developed March 18, 2013; alert roster developed April 10, 2013. COMMUNICATION PLAN HWY 70 REALIGNMENT As of 2/25/20 52 Assumption Parish Scientific Situation Summary University of Memphis:

As of 2/25/20 There were 112+ MEQs yesterday and 72 MEQs detected since midnight that meet the selection criteria. There were 4 VLP events detected yesterday and 2 VLPs detected since midnight. (12/26) There were 232 MEQs yesterday and 32+ MEQs detected since midnight that meet the selection criteria. There were 6 VLP events detected yesterday and no VLPs detected since midnight. (1/2/14) There were 2 MEQs yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (1/7/14) There was 1 MEQ yesterday and 4 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and 1 VLP detected since midnight. (1/10/14) There were 2 MEQs yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (1/14/14) There were 3 MEQs yesterday and 3 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (1/17/14) There were 2 MEQs yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (1/21/14) There were no MEQs yesterday and 4 MEQs detected since midnight that meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (1/24/14) There were 6 MEQs yesterday and 4 MEQs detected since midnight that meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (1/28/14) There were 3 MEQs yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (1/31/14) There were 2 MEQs yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (2/4/14) There were no MEQs yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (2/7/14) There was 1 MEQ yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight (2/11/14) There were 3 MEQs yesterday and 3 MEQs detected since midnight that meet the selection criteria. There was 1 VLP event detected yesterday and no VLPs detected since midnight. (2/14/14) There were 3 MEQs yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. The rate of MEQs is consistent with code 1, green. (2/18/14) 53 Assumption Parish Scientific Situation Summary University of Memphis: - - There were 3 MEQs yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (2/21/14) There were 3 MEQs yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (2/28/14) There were 6 MEQs yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLP detected since midnight. (3/7/14)

There were 2 MEQs yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLP detected since midnight. (3/11/14) There were 5 MEQs yesterday and no MEQs detected since midnight that meet the selection criteria. There was no VLP events detected yesterday and no VLP detected since midnight.(3/14/14) There was 1 MEQ yesterday and no MEQs detected since midnight that meet the selection criteria. There was no VLP events detected yesterday and no VLP detected since midnight. (3/18/14) There were 2 MEQs yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (3/21/14) There were 96 MEQs yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were 6 VLP events detected yesterday and no VLPs detected since midnight. Yesterday the Oxy 3 cavern pressure was bled down in the morning. A couple of hours later an intense episode of MEQ activity combined with large VLP events occurred. The seismic activity has significantly decreased overnight with no VLP events. (3/25/14) There were 2 MEQ yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (3/28/14) There was 1 MEQ yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (3/31/14) There was 1 MEQ yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/2/14) There was 1 MEQ yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/4/14) There were 4 MEQs yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/7/14) There was 1 MEQ yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/9/14) As of 2/25/20 54 Assumption Parish Scientific Situation Summary - University of Memphis: Dr Horton has observed 2 MEQ's since midnight and no VLP's (4/11/14) There were 4 MEQs yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/14/14) There was 1 MEQ yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/16/14) There were no MEQs yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/18/14) There were 5 MEQs yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/21/14) There was 1 MEQ yesterday and 1 MEQ detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/23/14) There was 1 MEQ yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/25/14) There were no MEQs yesterday and no MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/28/14) There was 1 MEQ yesterday and 4 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs detected since midnight. (4/30/14) There was 1 MEQ yesterday and 2 MEQs detected since midnight that meet the selection criteria. There were no VLP events detected yesterday and no VLPs

detected since midnight. (5/2/14) There were 6 MEQs yesterday and 2 MEQs since midnight that the meet the selection criteria. There was 1 VLP event detected yesterday and no VLPs detected since midnight. (5/5/14) There were 3 MEQs yesterday and no MEQs since midnight that the meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (5/7/14) There were 5 MEQs yesterday and 1 MEQ since midnight that the meet the selection criteria. There were no VLPs event detected yesterday and 1 VLP detected since midnight. (5/9/14) There were 6 MEQs yesterday and 3 MEQs since midnight that the meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (5/12/14) There were 6 MEQs yesterday and no MEQs since midnight that the meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (5/14/14) There were 2 MEQs yesterday and no MEQs since midnight that the meet the selection criteria. There were no VLPs event detected yesterday and no VLPs detected since midnight. (5/16/14) 55 As of 2/25/20 Assumption Parish Scientific Situation Summary University of Memphis and Magnitude: - There were 4 MEQs yesterday and no MEQs since midnight that the meet the selection criteria. There were 2 VLPs event detected yesterday and no VLPs detected since midnight. There were no MEQs yesterday and 1 MEQ since midnight that the meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (5/21/14) There was 1 MEQ yesterday and no MEQs since midnight that the meet the selection criteria. There were no VLPs? detected yesterday and no VLPs detected since midnight. (5/23/14) There were no MEQs yesterday and 1 MEQ since midnight that the meet the selection criteria. There was 1 VLP detected yesterday and no VLPs detected since midnight. (5/28/14) There were 5 MEQs yesterday and no MEQs since midnight that the meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (5/30/14) There were 3 MEQs yesterday and 1 MEQ since midnight that the meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (6/4/14). Magnitude notification system is now on line and reporting data and indicates low seismic activity. There were 2 MEQs yesterday and no MEQs since midnight that the meet the selection criteria. There were no VLP?s detected yesterday and 1 VLP detected since midnight. (6/11/14). Magnitude notification system is now on line and reporting data and indicates low seismic activity. There were 8 MEQs yesterday and 5 MEQs since midnight that the meet the selection criteria. There were no VLP detected yesterday and no VLPs detected since midnight. Magnitude notification system is now on line and reporting data and indicates low seismic activity. (6/25/14) There were 15 MEQs yesterday and 1 MEQ since midnight that the meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. The Magnitude system is nor on line and reporting data and indicates low seismic activity (7/2/14) There were 7 MEQs yesterday and 17 MEQs since midnight that the meet the selection criteria. There was 1 VLP detected yesterday and 2 VLPs detected since midnight. The Magnitude system is now on line and reporting data and indicates low seismic activity (7/9/14) There were 9 MEQs yesterday and 2 MEQs since midnight that the meet the selection criteria. There was 1 VLPs detected yesterday and no VLPs detected since midnight. The Magnitude system is now on line and reporting data and indicates low seismic activity (7/16/14) There were 12 MEQs yesterday and no MEQs since midnight that the meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. The Magnitude system is now on line and reporting data and indicates low seismic activity (7/23/14) There were 51+ (I stopped counting at 51. There were several hundred.) MEQs yesterday and 23 MEQs since midnight that meet the selection criteria. There were 4 VLPs detected yesterday and 11 VLPs detected since midnight. As seen on LA12 sustained VLP activity occurred for an hour or so after 2200 CDT. (8/6/14) There were 2 MEQs yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and 1 VLP detected since midnight. (8/13/14) There was 1 MEQ yesterday and 1 MEQ since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (8/20/14)

As of 2/25/20 56 Assumption Parish Scientific Situation Summary University of Memphis and Magnitude: - There were 3 MEQs yesterday and 4 MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. There were 5 MEQs yesterday and no MEQs since midnight that meet the selection criteria. There was 1 VLP detected yesterday and no VLPs detected since midnight. (9/3/14) There were 23 MEQs yesterday and 1 MEQ since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (9/10/14) There were 2 MEQs yesterday and 2 MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight (9/17/14) There were no MEQs yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (9/24/14) There were 4 MEQs yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (10/1/14) There were no MEQs yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (10/8/14) There were 2 MEQs yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (10/15/14) There was 1 MEQ yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (10/22/14) There were 2 MEQs yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (10/29/14) There were 3 MEQs yesterday and no MEQs since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (11/5/14) There were 2 MEQs yesterday and 1 MEQ since midnight that meet the selection criteria. There were no VLPs detected yesterday and no VLPs detected since midnight. (11/12/14) As of 2/25/20 57 Assumption Parish Scientific Situation Summary As of 2/25/20 Scientific Workgroup theories of cause: Salt Dome moving natural migration of gas Failed cement casing in OXY #3 well Cavity Failure Salt / Caprock falling from top of the cavern Natural Man-made (including penetration into sediments by cavern) Gas storage cavern connections, communications by fractures

Low permeability seepage of gas into OXY #3 (source unknown) Regional Tectonic activity (movement on growth faults) A combination of above events Held monthly meeting (conference call) on 2/19/14 58 Next Operational Period (12 - 19 Nov 14) Incident Action Plan The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing operational response activities. DNR: As of 2/25/20 Texas Brine conducting hydrocarbon removal operations Perform oversight of operations at investigatory well, observation/vent well and sink hole (day light hours) Staff providing assist Assumption Parish Incident Commander and DEQ with bubble site monitoring, community air monitoring, and in-home monitoring Reviewing analysis of water samples from water wells screened in Mississippi River Alluvial aquifer Continue source investigation and file reviews Coordinate and provide information to science team Dome operators proceeding to implement their Shallow Geology, Aquifer, and Caprock Characterization Plan CB&I to collect pressure readings from the shallow water Geoprobes CB&I continues to evaluate the production of formation gas from the vent wells November 12, 2012 emergency order from the Commissioner of Conservation: Homes Ventilation improvements in unventilated, unoccupied, enclosed spaces Gas Monitors Texas Brines contractor is Installing pressure monitoring wells with DNR oversight. Develop plan for mitigation problem over area Contain dissolved contamination from flowing into Bayou Corne 59 Next Operational Period (12 - 19 Nov 14) Incident Action Plan The State Agency is responsible for oversight of the following activities. Texas Brine and its contractors are performing

operational response activities. DNR (Contd) Texas Brine contractor to maintain antennas & repeaters for in-home monitoring Texas Brine contractor to maintain in-home monitors. BRC continues to evaluate data and provide recommended requirements Monitoring under slab venting installation Weekly surveys Conduct daily well readings and flare maintenance Monitor Bubble site 56, 5, 9, 23 Monitor continuous dewatering of ORWs PMWs - 04 and 07 - monitor for safety and security Monthly monitoring and maintenance of PVWs and KGCs Redevelop ORWs 6, 19, 37 Collect gas samples from PVWs and Flares Perform surface completions for MRAA wells 6D and 9D As of 2/25/20 60 Next Operational Period (12 - 19 Nov 14) Incident Action Plan DEQ: As of 2/25/20 Continue to conduct air and water quality monitoring in the community and in the area of the bubbling in Bayou Corne, Triche Canal, and Grand Bayou

Continue to offer indoor air monitoring to residents who complete the Rite of Passage form from the Parish. Residents desiring indoor home monitoring may call DEQ @ (225) 2193015 Pursuant to the DNR press release dated 11/7/2012 regarding the pressure in some geoprobes, LDEQ will conduct indoor air monitoring as requested at residential homes in the Bayou Corne community. Conduct any test required by scientific group to aid in source/cause identification Provide oversight of Texas Brines efforts to conduct hydrocarbon removal at the slurry site. Provide oversight and review of Texas Brines air and water quality monitoring efforts. Continue to visually monitor conditions of the booms and general conditions of the slurry area from the well pad. Perform Bayou Corne water quality background sampling Public hearing on September 9th at 6:00 pm 61 Next Operational Period (12 - 19 Nov 14) Incident Action Plan DHH / Office of Public Health DHH will continue to review and analyze environmental sample data as provided by the Louisiana Department of Environmental Quality (LDEQ), for any chemical findings that would indicate a potential public health risk. DHH Safe Drinking Water Program will continue to monitor the incident for any reports of the expansion of the slurry/hole to ensure that water distribution underground piping is not compromised. As of 2/25/20 62 Next Operational Period (12 - 19 Nov 14) Incident Action Plan DHH / Office of Public Health: DHH/OPH will continue to participate in meetings with GOHSEP and Assumption Parish to maintain visibility and operational readiness for any potential public health concerns. As of 2/25/20 63 Next Operational Period (12 - 19 Nov 14) Incident Action Plan DOTD:

As of 2/25/20 Continue daily observation of road and bridges Will continue observation and analysis of deformation spots along Hwy 70 in evacuation zone Continuous monitoring of Hwy 70 with instrumentation Continue observation and analysis of Hwy 69 Continue establishing a fifth CORS site in the Bayou Corne area 64 Next Operational Period (12 - 19 Nov 14) Incident Action Plan Scientific Group: Analyzing groundwater sample analyses Evaluate issues as they arise Continue to provide information to the UCG Anyone can submit information at any time for science team evaluation by sending an email to [email protected] Additional members added to the team as needed As of 2/25/20 65 Assumption Parish Long Range Plan DNR: Conducting testing of investigatory well to determine cavern condition and root cause Scientific Group conference calls/meetings will be scheduled weekly to review new data and issues that arise Continue source investigation and file review

Evaluate results of monitoring data and take necessary action Monitor operations at investigatory, observation/vent well and sink hole (day light hours) Staff on standby to assist incident commander Ensure total remediation of the entire site Continue to pursue landowner access agreements Implement CB&Is Plan of Action Conduct industrial well sampling bi-weekly and bubble sites as reported in general vicinity of operations. Dome operators to identify presence of natural gas and vent as necessary Ensure Texas Brine fully complies with the Commissioner of Conservations Orders dated 11 Oct 12, Nov 2012, 7 Dec 2012, and 14 Jan 2013. Blue Ribbon Commission will evaluate data collected in order to provide recommendations on risk levels to the Parish DOTD: Investigating bypass alignment of LA 70 around the Salt Dome Continuous automated monitoring of roadway-bridge movement and subsidence Establish CORS site #5 on the North side of Hwy 70 DEQ: Continue to conduct sampling of air and water in area of operation Continue oversight of ongoing remedial activities Continue to conduct hazard assessment monitoring with APSO in resident homes and properties when requested. As of 2/25/20 66

Recently Viewed Presentations

  • Energía - Blog de Física y Química | IES Juan de la Cierva

    Energía - Blog de Física y Química | IES Juan de la Cierva

    … entonces cuando el cuerpo esté a una altura h podemos decir que la energía potencial gravitatoria es: Ep = m g h Trabajo y variación de energía potencial gravitatoria suelo Ep = 0 cuando h = 0 h Ep...
  • Programming Languages & Software Engineering

    Programming Languages & Software Engineering

    fork. method of . RecursiveAction. calls . compute() in a new thread/task. Calling . compute. directly is a plain-old method call. join. method of . RecursiveAction. blocks its caller until/unless the receiver is done executing (its . compute. returns) Must....
  • HCV and IR

    HCV and IR

    Excess extrahepatic mortality associated with HCVThe REVEAL HCV Cohort Study. 19,636 HBsAg-seronegative adults, aged 30-65 yrs. 1,095 anti-HCV+ [5.6%] 2,394 deaths after an average FU of 16.2 years
  • The Colorado Academic Standards

    The Colorado Academic Standards

    Consider released items for a PBA set at one grade level: What comprehension and critical thinking strategies will students need to utilize to be successful on this PBA? How are the standards are being assessed? Participants examine sample test items...
  • Everything You Need to Know about SMTP Transport for Office 365

    Everything You Need to Know about SMTP Transport for Office 365

    Everything You Need To Know about SMTP Transport for Microsoft Office 365. Brian Reid - C7 Solutions Ltd. ... Office 365 mailbox > Exchange Online Protection > MX for target domain ... Everything You Need to Know about SMTP Transport...
  • Plant-feeding mite pests - entnemdept.ufl.edu

    Plant-feeding mite pests - entnemdept.ufl.edu

    Plant-feeding mite pests. Photo credit: Gary R. Bauchan, USDA-ARS Electron & Confocal Microscopy Unit. SEM of adult red palm mite, Raoiella indica In this presentation we will discuss the basics of mite biology and anatomy and learn some of the...
  • eRA Commons: Accounts, Roles, Profiles & Delegations

    eRA Commons: Accounts, Roles, Profiles & Delegations

    Account - Tips. Designate more than one Signing Official. Certain Commons actions can only be taken by users with the SO role and you will want a back-up. Administrators should always check to see if a user already has a...
  • New Laws 2018 - adi-sandiego.com

    New Laws 2018 - adi-sandiego.com

    A BUMPER CROP! There's a larger number of important new criminal law bills this year than any since I started compiling these in 1999. We've had years of greater overall importance, like 2011's Realignment's largest change in crim. law since...